### v1

# Philip ${}^{*\,\sharp}$ Yandex, MIPT &Aleksandr Beznosikov

Abstract
With the increase in the number of parameters in large language models, the process of pre-training and fine-tuning increasingly demands larger volumes of GPU memory. A significant portion of this memory is typically consumed by the optimizer state. To overcome this challenge, recent approaches such as low-rank adaptation (LoRA (Hu et al., 2021)), low-rank gradient projection (GaLore (Zhao et al., 2024a)), and blockwise optimization (BAdam (Luo et al., 2024)) have been proposed. However, in all these algorithms, the effective rank of the weight updates remains low-rank, which can lead to a substantial loss of information from the gradient. This loss can be critically important, especially during the pre-training stage. In this paper, we introduce FRUGAL (Full-Rank Updates with GrAdient spLitting), a new memory-efficient optimization framework. FRUGAL leverages gradient splitting to perform low-dimensional updates using advanced algorithms (such as Adam), while updates along the remaining directions are executed via state-free methods like SGD or signSGD (Bernstein et al., 2018). Our framework can be integrated with various low-rank update selection techniques, including GaLore and BAdam. We provide theoretical convergence guarantees for our framework when using SGDM for low-dimensional updates and SGD for state-free updates. Additionally, our method consistently outperforms concurrent approaches across various fixed memory budgets, achieving state-of-the-art results in pre-training and fine-tuning tasks while balancing memory efficiency and performance metrics.
\$*\$\$*\$footnotetext: Work was mostly done during the internship at Mohamed bin Zayed University of Artificial Intelligence.$\sharp$\$$\sharp$\$footnotetext: The corresponding author: zmushko.ph.a@gmail.com\$$\diamond$\$$\diamond$\$footnotetext: Moscow Institute of Physics and Technology §$\S$ Mohamed bin Zayed University of Artificial Intelligence\$$\dagger$\$$\dagger$\$footnotetext: Ivannikov Institute for System Programming of the Russian Academy of Sciences $\ddagger$ Innopolis University
1 Introduction
In recent years, Large Language Models (LLMs) such as GPT (OpenAI, 2023) and LLaMA-3 Dubey et al. (2024) have demonstrated remarkable performance across various disciplines (Brown, 2020; Yang et al., 2024; Romera-Paredes et al., 2024). However, a critical factor in achieving these results is the size of these models (Hoffmann et al., 2022). A larger number of parameters not only increases computational cost but also significantly raises memory requirements. For instance, training an 8 billion parameter LLaMA model in a 16-bit format necessitates each parameter to occupy 2 bytes, resulting in 16GB for storing the parameters and an additional 16GB for gradients. Utilizing the Adam optimizer (Kingma, 2014), which is standard for pre-training and fine-tuning LLMs, adds a further 32GB of memory to store the $m$ and $v$ statistics, resulting in 64GB total amount of memory. Furthermore, to achieve higher-quality results, training in pure 16-bit format is often insufficient (Zamirai et al., 2020). This necessitates storing master weights and optimizer statistics in 32-bit format, leading to total memory demands that exceed the capacity of cutting-edge graphics cards, such as the A100-80GB.
Numerous research projects have been aimed at reducing these significant costs. These approaches include engineering solutions like gradient checkpointing Chen et al. (2016) and memory offloading (Rajbhandari et al., 2020), which do not change the training trajectory. There are also methods that adjust the training algorithm by decreasing the number of trainable parameters (Frankle & Carbin, 2018; Wang et al., 2023; Sreenivasan et al., 2022; Horváth et al., 2024) or their bit precision (Wortsman et al., 2023), as well as optimizer statistics (Dettmers et al., 2021; Shazeer & Stern, 2018; Zhang et al., 2024c). In this work, we concentrate on the latter category.
Parameter-Efficient Fine-Tuning (PEFT) methods, such as LoRA (Hu et al., 2021), Dora (Liu et al., 2024), and BitFit (Zaken et al., 2021) reduce memory costs by training a relatively small number of parameters compared to the size of the original model, while the remaining modules are frozen. This approach has proven effective for the task of efficient fine-tuning of pre-trained language models. However, PEFT methods have a fundamental limitation: parameter updates always lie in a low-dimensional subspace, which prevents the use of these methods for pre-training (Lialin et al., 2023) and may constrain their capabilities in fine-tuning (Zhang et al., 2024a).
Recent works, such as GaLore (Zhao et al., 2024a), ReLoRA (Lialin et al., 2023), and BAdam (Luo et al., 2024), offer a solution to this problem. These methods enable higher-dimensional full-parameter learning by periodically changing the optimizable low-rank subspace, which we will call the state-full subspace. However, even though these methods result in overall parameter changes that are high-dimensional, the updates in each step remain low-dimensional. The dimensionality of the state-free subspace significantly exceeds that of the state-full subspace. The remaining information contained in the gradient is not utilized for parameter updates. Nevertheless, this information can still be leveraged to train the model.
Figure 1: FRUGAL reduces memory usage by splitting gradient updates into low-dimensional updates with advanced optimizers (e.g., AdamW) and using state-free methods (e.g., SignSGD and SGD) for the rest.
Algorithm 1 FRUGAL(State-Full, State-Free)
Input: model $f_{\mathbf{\theta}}(\cdot)$ with $p$ parameters $\{\theta_{i}\in\mathbb{R}^{d_{i}}\}_{i=1}^{p}$, loss function $\mathcal{L}$, gradient projectors $P_{k,i}$ for $i\in[p],k\in[K]$.
1:for $k=1,2,\ldots K$ do
2:     get data batch $(x,y)$
3:     compute $\ell\leftarrow\mathcal{L}(f_{\theta}(x),y)$triangleright Forward
4:     for $g_{i}=\frac{\partial\ell}{\partial\theta_{i}}$ from Backward do
5:         $g_{\text{full},i}\leftarrow P_{k,i}(g_{i}),\;$triangleright Project Grad
6:         $g_{\text{free, i}}\leftarrow g_{i}-P^{-1}_{k,i}(g_{\text{full},i})$triangleright Residual
7:         $\texttt{s}_{\theta_{i}}\leftarrow[P_{k,i}(P^{-1}_{k-1,i}(s),\;s\in\texttt{s}_{\theta_{i}}]$$\triangleright$ Project States
8:         $u_{\text{full, i}}\leftarrow\texttt{State-Full.update}(\theta_{i},g_{\text{full},i},\texttt{s}_{\theta_{i}})$
9:         $u_{\text{free, i}}\leftarrow\texttt{State-Free.update}(\theta_{i},g_{\text{free},i})$
10:         $\theta_{i}\leftarrow\theta_{i}+P^{-1}_{k,i}(u_{\text{full},i})+u_{\text{free, i}}$
11:     end for
12:end for
We present the FRUGAL framework, designed to bridge this gap. Our approach stems from a crucial observation: although memory constraints prevent the use of advanced optimizers like Adam (Kingma, 2014) in the state-free subspace, one still can update this subspace using state-free optimization algorithms like Stochastic Gradient Descent (SGD) or signSGD (Bernstein et al., 2018). This solution allows for high-dimensional updates, which provides additional opportunities to explore the parameter space and improves convergence.
Contributions. We summarize the main contributions of our work as follows:
  • We present a new memory-efficient optimization framework that combines the use of advanced optimization algorithms for the state-full subspace with state-free algorithms for the complementary subspace. The framework supports various types of state-full optimizers, state-free optimizers, and different methods for projecting the gradient onto the state-full subspace.
  • We provide theoretical convergence guarantees for our framework. In the proof, we consider the case where SGDM acts as the state-full optimizer and SGD as the state-free optimizer, and we show that FRUGAL matches the best-known convergence rate in many scenarios.
  • To verify the practical applicability of FRUGAL, we conduct extensive experiments in popular real-world scenarios. [^1] The code is available at https://github.com/fzmushko/FRUGAL. In these experiments, we pre-train LLaMA-like models (up to 1B parameters) on the Colossal Clean Crawled Corpus (C4) dataset (Raffel et al., 2020) and fine-tune RoBERTa (Liu, 2019) on the GLUE benchmark (Wang, 2018). The results show that our method significantly outperforms previous memory-efficient algorithms while using the same memory budget.
  • We demonstrate that only the Logits layer in transformer-like models requires advanced optimizers like Adam, while other modules (including Embeddings and RMSNorms) can use simpler methods like signSGD without significant performance loss. This opens new possibilities for memory-efficient training and provides crucial insights into the learning dynamics of Transformers.
2 Related work
Memory-efficient full-parameter learning. Recent research has focused on reducing the memory footprint of LLMs by decreasing the size of optimizer states while maintaining their performance. Low-rank adaptation methods, such as LoRA (Hu et al., 2021), inject trainable rank decomposition matrices into each layer of the model, reducing memory requirements by optimizing only a few learnable adapters. ReLora (Lialin et al., 2023) builds upon this by merging low-rank adaptations into the main model weights during training, potentially increasing the total rank of the update. BAdam (Luo et al., 2024) leverages Block Coordinate Descent for full-parameter training by switching active blocks during fine-tuning. MicroAdam (Modoranu et al., 2024) compresses gradient information before feeding it into the optimizer state, significantly reducing memory footprint while enabling full parameter learning through error feedback mechanisms. GaLore (Zhao et al., 2024a) maintains full parameter learning by projecting gradients onto a low-rank subspace using truncated SVD decomposition, storing optimizer states in this reduced space. Notably, GaLore achieves good results in pre-training, with performance close to that of Adam. However, while these methods effectively reduce memory overhead, they all perform low-rank updates at each iteration. In contrast, our approach utilizes all available gradient information to perform full-dimensional updates at each optimizer step, offering a novel perspective on memory-efficient optimization for LLMs.
Other memory-efficient optimization. Several other methods have been proposed to reduce the memory footprint of optimizers. AdaFactor (Shazeer & Stern, 2018) attempts to mimic Adam’s behavior while reducing memory usage through factorization of the variance matrix $v$. Adam-mini (Zhang et al., 2024c) further reduces memory by storing only one value $v$ per block. Dettmers et al. (2021) and Li et al. (2024) decrease memory footprint by quantizing optimizer states to the lower-precision representations. Lv et al. (2023) proposed to reduce weight gradient memory by fusing the backward operation with the optimizer update. Notably, these approaches are orthogonal to our method FRUGAL and can be combined with it for further memory efficiency.
Block Coordinate Descent. Block Coordinate Descent (BCD) is a well-established optimization method with a rich history in mathematical optimization (Ortega & Rheinboldt, 2000; Tseng, 2001; Richtárik & Takáč, 2014; 2015b; Richtárik & Takác, 2016; Takáč et al., 2013; Richtárik & Takáč, 2015a). In recent years, a specific instance of BCD, known as layer-wise learning, has been applied to deep learning. Notable examples include (Luo et al., 2024; Pan et al., 2024), which leverage this approach for LLM fine-tuning. To the best of our knowledge, our work presents the first theoretical analysis of an extended BCD framework (Section 4) where the remaining coordinates are also updated using a different algorithm. This novel approach extends traditional BCD techniques, opening new avenues for full model optimization in deep learning.
Sign-based methods for training language models. Since its introduction, Adam has become the de facto primary optimization algorithm, demonstrating superior practical results compared to SGD-based algorithms across various deep learning tasks. This difference is particularly noticeable when training Transformers on language tasks. While Zhang et al. (2020) hypothesized that Adam outperforms SGD in this setup due to the heavy-tailed distribution of sampling-induced errors, Kunstner et al. (2023) demonstrated that this superiority persists even in full-batch training. They proposed a new hypothesis suggesting that Adam’s key success factor is related to its similarity to signSGD (Balles & Hennig, 2018; Balles et al., 2020), and both Kunstner et al. (2023) and Zhao et al. (2024b) showed that signed descent with momentum reduces the performance gap with Adam. In contrast, to the best of our knowledge, we are the first to train the majority of language model’s parameters using signSGD without momentum, achieving minimal loss in quality. This approach further demonstrates the effectiveness of sign-based methods for LLM training, paving the way for more efficient and scalable optimization strategies in deep learning.
Table 1: Comparison of different projection and optimization strategies on pre-training LLaMA-130M on C4 with Adam as the state-full algorithm. State-free subspace is updated for full-rank methods but frozen for low-rank methods. Low-rank SVD outperforms Low-rank Random at the beginning of training but falls behind later due to insufficient exploration of the parameter space.
Method Validation perplexity after iterations $\downarrow$
4k 20k 40k 100k 200k
Low-rank SVD 39.75 26.35 24.38 22.30 21.11
Low-rank Random 42.31 25.99 23.55 21.33 20.01
Full-rank Random 37.26 23.46 21.53 19.66 18.64
Full-rank SVD 33.96 22.54 21.01 19.30 18.35
Full-rank RandK 36.38 23.02 21.25 19.70 18.63
Full-rank Block 37.20 23.34 21.42 19.59 18.60
Adam 33.95 21.90 20.56 18.97 18.13
3 FRUGAL: Full-Rank Updates with GrAdient spLitting
In this section, we first introduce two main building blocks of our method and then introduce our general framework.
3.1 The importance of exploring the entire space during the training process
In recent work, Zhao et al. (2024a) proposed GaLore, an optimization method based on projecting the gradient matrix ${\bm{G}}$ of each layer onto a low-dimensional subspace. To obtain the projection matrix ${\bm{P}}$, they use SVD decomposition of ${\bm{G}}_{t}$, which is recomputed with frequency $T$. The vectors or rows of ${\bm{G}}$ are projected onto the first $r$ left or right singular vectors, respectively. This approach has theoretical foundations: the first $r$ singular vectors correspond to the first $r$ singular values and, therefore, should better utilize information from the spectrum of ${\bm{G}}$.
The authors pointed out that calculating the SVD decomposition results in extra computational overhead, which can be as much as a 10% increase as the hidden size of the model grows. To minimize this cost and examine the significance of using SVD decomposition, one may wonder about the possibility of employing a random semi-orthogonal projection matrix ${\bm{R}}$ instead of projecting onto the first $r$ singular columns with ${\bm{P}}$. Surprisingly, we found that although SVD decomposition delivers an initial boost, subsequent training with random projection yields significant improvements. As an illustration, we took the pre-training [^2] See Section 5.1 for a detailed description and discussion on the experimental setup. of a 130M model with LLaMA-like architecture on the C4 dataset (Raffel et al., 2020). The results are presented in Table 1, where training with ${\bm{P}}$ corresponds to the row labeled “Low-rank SVD”, and training with ${\bm{R}}$ corresponds to the row labeled “Low-rank Random”. The ranks of both projections ${\bm{P}}$ and ${\bm{R}}$ equal $192$.
To investigate this phenomenon, we pre-trained the LLaMA-60M model and collected gradients ${\bm{G}}_{t}$ from different iterations $t$ for examination. Following the setup from GaLore (Zhao et al., 2024a), we computed SVD decompositions and extracted projections ${\bm{P}}_{t}$ with a rank of 128. We evaluated the similarity of the projection matrices by calculating the principal angles between different projections ${\bm{P}}_{t}$ at different steps. Similarly to the observations in Q-Galore (Zhang et al., 2024d), we found that these projections show minimal change during the training period; see Figure 2.
Here, we take the projection matrix corresponding to the $5$-th layer and plot histograms of the cosine of the principal angles between pairs ${\bm{P}}_{t}$ and ${\bm{P}}_{t^{\prime}}$ from different iterations. For comparison, we also include the random projections on the right. As can be seen, the distributions of cosines differ significantly for the ${\bm{P}}_{t}$ and for the random projections. While ${\bm{R}}_{t}$ feature no angles with cosines higher than $0.9,$ the top $57$ cosines for ${\bm{P}}_{t}$ surpass $0.9$, even for gradients $1000$ steps apart.
This leads to the conclusion that although SVD decomposition generally better captures the information contained in the ${\bm{G}}_{t}$, the original GaLore algorithm updates weights only in a small subspace. We hypothesize that training with random projections yields superior results due to the more extensive investigation of the optimizable space during the training process. This finding indicates that to achieve better convergence, it is important to seek out optimization algorithms that explore the entire space during the training process.
3.2 Advantage of the Full-Rank Updates
The insight from the Section 3.1 suggests that the training of language models performs significantly better when the entire parameter space is utilized during the training process. Given the importance of updating parameters in all directions, this poses the question: Is it optimal to use low-rank updates, as employed by methods such as GaLore (Zhao et al., 2024a), ReLoRA (Lialin et al., 2023), and BAdam (Luo et al., 2024)? Using low-rank updates means the effective rank of the update is significantly smaller than the full dimensionality of the parameter space, inevitably leading to a loss of valuable information contained in the gradient.
However, the method to leverage the full-rank gradient for updating parameters is not readily obvious. Using algorithms like Adam (Kingma, 2014) is not an option due to the memory overhead they introduce, which is precisely what we aim to avoid. An alternative approach is to use state-free optimizers such as SGD or signSGD (Bernstein et al., 2018). Unfortunately, SGD have been shown to be ineffective for training transformer models, as shown in Zhang et al. (2020); Pan & Li (2023).
Figure 2: Histograms of principal angle cosines. The first three are taken between ${\bm{P}}_{t}$ and ${\bm{P}}_{t^{\prime}}$ from different iterations $t$ and $t^{\prime}$. ${\bm{P}}$ is obtained from the truncated SVD decomposition of the gradient ${\bm{G}}$ of the Key projection from the 5th layer. The last histogram is taken between two random semi-orthogonal projections ${\bm{R}}$ and ${\bm{R}}^{\prime}$ for comparison.
Nevertheless, a recent study Zhao et al. (2024b) suggests a promising methodology: while SGDM doesn’t generally work well with transformers, using SGDM for the majority of parameters and Adam for a selected subset can lead to effective training. This raises the question: could a hybrid approach using SGD or signSGD instead of SGDM be viable? If the key subset of parameters is handled by advanced algorithms, can the other parameters be trained effectively with state-free optimizers?
To address this question, we conducted an experiment on LLaMA-130m, where we utilized the Adam (Kingma, 2014) for state-full parameters and signSGD (Bernstein et al., 2018) for state-free parameters. A detailed description of the experimental setup can be found at Section A.1. The result is highlighted as the entry “Full-rank Random” in Table 1. Full-rank updates significantly enhance performance, approaching the efficiency of the memory-intensive Adam optimizer, which serves as a upper bound in terms of performance. These findings underscore the potential of state-free algorithms for updating a substantial portion of the parameter space, paving the way for efficient, scalable optimization methods that deliver high performance without the significant memory costs traditionally associated with state-of-the-art optimizers.
3.3 FRUGAL as Optimization Framework
The setup outlined at the conclusion of the Section 3.2 results in a general framework for memory-efficient optimization. Briefly, it operates as follows: the entire space is partitioned into state-full and state-free subspaces. The state-full subspace is updated using advanced algorithms, while the state-free subspace is updated using a state-free method. After a certain number of steps, the state-full subspace is changed to better explore the optimization space. However, determining the optimal state-free optimizer and the projection method onto the state-full subspace is not readily apparent. In this section, we strive to find the optimal configuration.
State-free optimizer. We conducted a preliminary experiment updating all parameters using state-free algorithms to choose between SGD and signSGD (Bernstein et al., 2018). Table 8 presents these results. After testing various learning rates, we found that signSGD consistently outperforms SGD, leading us to favor signSGD. We attribute this performance to the similarities between signSGD and Adam (Kingma, 2014), as noted in Balles & Hennig (2018); Balles et al. (2020); Kunstner et al. (2023). Additionally, signSGD produces updates of similar magnitude to those generated by Adam, which simplifies the calibration of the learning rate for state-free parameters.
Projection type. When selecting a projection method, it is crucial to strike a balance between quality and memory efficiency. For instance, the approach using the projection matrix from SVD decomposition, as in GaLore (Zhao et al., 2024a), better preserves the information embedded in the gradient but requires additional memory for storing projection matrices and computational resources for performing the SVD. To reduce computational demands, one could employ random coordinate projection denoted as RandK, but this requires additional memory or recomputation [^3] See Appendix B for discussion on the memory requirements for different projection methods.. A more structured alternative is to select not random matrix elements but entire random columns or rows. The most aggressive approach follows the method from BAdam, wherein an entire block is chosen as the state-full subspace.
The performance results obtained with all these variants are presented in Table 1. Full-rank SVD outperforms both Full-rank RandK and Full-rank Block methods, demonstrating comparable performance. The superior performance of SVD projections can be explained by their ability to extract the principal information from the gradient. Nonetheless, a downside is the increased compute and memory demand from SVD. Therefore, we opt for the blockwise selection, as it is the most memory-efficient — requiring only the storage of active block indices.
A formal description of the final algorithm is presented in Algorithm 1. It’s worth noting that this framework allows for variation not only in the State-Full algorithm but also in the choice of projection and State-Free optimizer. In our experiments in Section 5, we use a specific variant with Adam as the State-Full optimizer, signSGD as the State-Free optimizer, and blockwise projection.
For Line 7, state projection, in Algorithm 1, we note that if the projection does not change, i.e., $P_{k,i}=P_{k-1,i}$, then $P_{k,i}(P^{-1}_{k-1,i}(s))=s$. Thus, we only need to project states when the projection changes from one round to another. However, our preliminary experiments with RandK selection showed that resetting states performs comparably to projection. Therefore, we could replace this projection with state resetting when the projection changes, which also aligns with blockwise subspace selection. However, either resetting or projecting states is important since we want projected gradients and optimizer states to reside in the same space. For instance, GaLore ignores this step, which leads to degraded performance when projections are updated frequently; see Appendix C for details.
Algorithm 2 FRUGAL(SGDM, SGD)
Input: momentum weight $\beta\in[0,1)$, initialization $x^{1}\in\operatorname*{\mathbb{R}^{d}}$ and $m^{0}=0$, step sizes $\{\alpha^{k}>0\}_{k=1}^{K}$, momentum set $J_{k}\subset[d]$ for $k=1,2,\ldots$.
1:for $k=1,2,\ldots$ do
2:     Compute stochastic gradient $\tilde{g}^{k}\leftarrow\nabla f_{\zeta^{k}}(x^{k})$;
3:     Update momentum vector $\tilde{m}_{j}^{k}\leftarrow(1-\beta)\tilde{g}_{j}^{k}+\beta\begin{cases}\tilde{m}_{j}^{k-1}&\text{if }j\in J_{k},\\ 0&\text{otherwise;}\end{cases}$
4:     Compute update vector $\tilde{u}_{j}^{k}\leftarrow\begin{cases}\tilde{m}_{j}^{k}&\text{if }j\in J_{k},\\ \tilde{g}_{j}^{k}&\text{otherwise;}\end{cases}$
5:     Update iterate $x^{k+1}\leftarrow x^{k}-\alpha^{k}\tilde{u}^{k}$;
6:end for
4 Theoretical Results
For the theoretical analysis, we consider the case where the State-Free optimizer is SGD and the State-Full optimizer is SGD with momentum (SGDM). For the projection, we use coordinate-wise projection. This special case of FRUGAL is provided in Algorithm 2. We minimize the following objective
$\textstyle\min_{x\in\mathbb{R}^{d}}\big{\{}f(x):=\mathbb{E}_{\zeta}[f_{\zeta}(x)]\big{\}},$ (1)
where we access $f$ via a stochastic oracle that takes $x$ as input and returns $(f_{\zeta}(x),\nabla f_{\zeta}(x))$.
4.1 Notation and Preliminaries
[…]
Assumption 1.
We make the following assumptions, which are standard in non-convex stochastic optimization; see (Liu et al., 2020).
  1. 1.
    Smoothness: The objective $f(x)$ in equation 1 is $L$-smooth.
  2. 2.
    Unbiasedness: At each iteration $k$, $\tilde{g}^{k}$ satisfies $\mathbb{E}_{\zeta^{k}}[\tilde{g}^{k}]=g^{k}$.
  3. 3.
    Independent samples: The random samples $\{\zeta_{k}\}_{k=1}^{\infty}$ are independent.
  4. 4.
    Bounded variance: The variance of $\tilde{g}_{j}^{k}$ with respect to $\zeta^{k}$ satisfies $\mathrm{Var}_{\zeta^{k}}(\tilde{g}_{j}^{k})=\mathbb{E}_{\zeta^{k}}[\|\tilde{g}_{j}^{k}-g_{j}^{k}\|^{2}]\leq\sigma_{j}^{2}$ for some $\sigma_{j}^{2}>0$. We denote $\sigma^{2}=\sum_{j=1}^{d}\sigma^{2}_{j}.$
Finally, we define the probability that index $j\in J_{k}$ is selected, conditioned on the prior iteration $k-1$, as $p^{k}_{j}:=\Pr_{k-1}[j\in J_{k}].$ Other useful quantities are $p^{k}_{\max}:=\max_{j\in[d]}\{p^{k}_{j}\}$ and $p^{k}_{\min}:=\min_{j\in[d]}\{p^{k}_{j}\}$.
4.2 Convergence of Algorithm 2
Below, we present the main convergence theorem.
Theorem 1.
Let Assumption 1 hold and $\alpha^{k}=\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$. Then, the iterates of Algorithm 2 satisfy
$\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}[\|g^{i}\|^{2}]$ $=\mathcal{O}\left(\frac{f(x^{1})-f^{*}}{k\alpha}+L\alpha\sigma^{2}\left(1+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{(1-\beta)}\right)\right),$ (2)
where $\bar{p}^{k}_{\min}=\frac{1}{k}\sum_{i=1}^{k}\bar{p}^{i}_{\min}$ and $\hat{p}^{k}_{\max}=\max_{i\in[k]}\{p^{i}_{\max}\}$.
The proof is deferred to Appendix E. Let us analyze the obtained result. Firstly, if $J_{k}=[d]$ or $J_{k}=\emptyset$, Algorithm 2 becomes SGDM and SGD, respectively. In this case, we have $\bar{p}^{k}_{\min}=1$ for SGDM and $\hat{p}^{k}_{\max}=0$ for SGD. Therefore, the resulting rate is $\mathcal{O}\left(\nicefrac{{1}}{{k\alpha}}+L\alpha\sigma^{2}\right)$, which recovers the best-known rate for both SGD and SGDM under these assumptions Liu et al. (2020). Furthermore, if at each step each coordinate is sampled independently with probability $p$, we have $\bar{p}^{k}_{\min}=\hat{p}^{k}_{\max}=p$. Therefore, we recover the same rate if $p=\mathcal{O}\left(1-\beta\right)$ or $p=\mathcal{O}\left(\beta\right)$. Finally, in the worst case (e.g., $J_{k}$ is deterministic and $0<|J_{k}|<d$), we have $\bar{p}^{k}_{\min}=0$ and $\hat{p}^{k}_{\max}=1$. Thus, the rate becomes $\mathcal{O}\left(\nicefrac{{1}}{{k\alpha}}+\nicefrac{{L\alpha\sigma^{2}}}{{1-\beta}}\right)$, which is worse by a factor of $\nicefrac{{1}}{{1-\beta}}$. However, this is expected since the bias from momentum is not outweighed by the variance reduction effect, as only the coordinates with momentum enjoy reduced variance; see Lemmas 1 and 2 in the appendix for details.
5 Experiments
This section presents the main experimental results of the paper. To evaluate the performance of FRUGAL, we conducted experiments both on the pre-training and fine-tuning of language models.
5.1 Pre-training experiments
Setup. The core setup for pre-training is taken from the Zhao et al. (2024a). We utilize LLaMA-based (Touvron et al., 2023a) model architectures with up to 1B parameters and train them on the Colossal Clean Crawled Corpus (C4) dataset (Raffel et al., 2020). The C4 dataset is intended for pre-training, making this setup a good approximation of real-world applications. A detailed description of the setup can be found in Section A.1. However, we made several modifications that we would like to discuss in detail below.
  • Training Duration. The training approach in Zhao et al. (2024a) aligns with the empirical rule from scaling laws (Hoffmann et al., 2022), which suggests using approximately 20 times the model size in tokens for training. However, this number of tokens is far from achieving convergence. In practice, models are typically trained for significantly longer periods (Touvron et al., 2023b; Zhang et al., 2024b). One reason for this discrepancy is that the original scaling laws do not account for the inference of the model after training. Adjustments to scaling laws considering this parameter are discussed, for example, in (Sardana & Frankle, 2023). For our experiments we chose 200k steps for the 60M and 130M models, 240k for 350M model and 300k for the 1B model.
  • Learning Rate. The authors of GaLore suggested using different learning rates for fixed unprojectable parameters (Embeddings, RMSNorms (Zhang & Sennrich, 2019), Logits) and the remaining projectable parameters (attention and FFN weights modules weights). However, introducing additional hyperparameters complicates the use of the algorithm. Since both sets of parameters are state-full and trained using the same optimization algorithm, we always used the same learning rate for them in FRUGAL and BAdam. For GaLore learning rate see Section 5.1.
  • Mixed Precision instead of the pure bfloat16 training. Pure 16-bit training has been shown to potentially compromise model convergence and accuracy (Zamirai et al., 2020). This degradation stems from storing both the model weights and optimizer statistics in reduced precision formats such as float16 or bfloat16. However, these formats often lack sufficient precision in representing floating-point numbers. Consequently, mixed precision training has become a more common approach for training language models (Le Scao et al., 2023; Almazrouei et al., 2023)). While training in pure 16-bit format is also possible, stochastic rounding (Gupta et al., 2015; Zamirai et al., 2020) is often employed to mitigate the aforementioned issue. Given that the goal of this research is to identify the optimal optimization algorithm, we deemed it more appropriate to compare optimizers in a transparent and stable setup that does not require auxiliary tricks. Hence, we primarily used Mixed Precision training for its illustrative value in understanding each method’s potential. However, for completeness, we also conducted experiments in pure bfloat16 format, detailed in our ablation study Section 5.1.2.
Table 2: Comparison of validation perplexity and memory estimation for various optimization methods across LLaMA model scales trained on C4. Lower $\downarrow$ is better. In parentheses, we also indicate the additional memory overhead introduced by the optimization algorithm. The values are calculated assuming that each float value occupies 4 bytes (float32). Note that Embeddings, RMSNorms, and Logits are always trained using state-full Adam.
60M 130M 350M 1B
Adam 22.73 (0.43G) 18.13 (1.00G) 14.43 (2.74G) 12.02 (9.98G)
GaLore, $\rho=0.25$ 25.68 (0.30G) 21.11 (0.54G) 16.88 (1.10G) 13.69 (3.41G)
BAdam, $\rho=0.25$ 24.86 (0.29G) 20.34 (0.52G) 16.41 (1.05G) 13.75 (3.23G)
FRUGAL, $\rho=0.25$ 23.59 (0.29G) 18.60 (0.52G) 14.79 (1.05G) 12.32 (3.23G)
FRUGAL, $\rho=0.0$ 24.06 (0.24G) 18.90 (0.37G) 15.03 (0.49G) 12.63 (0.98G)
Training tokens 20B 20B 24B 30B
Number of iterations 200k 200k 240k 300k
5.1.1 Comparison to existing memory-efficient algorithms
To begin, we present the results of comparing FRUGAL with existing memory-efficient methods across four sizes of LLaMA-based architectures: 60M, 130M, 350M, and 1B [^4] See preliminary experimental results with LLaMA 7B in Appendix D.
Baselines. We use the following methods as baselines for our approach:
  • Full-rank Training. Training using memory-inefficient Adam. Weights, gradients, and statistics are stored and computed for all parameters. This serves as an upper bound for model performance.
  • GaLore. Zhao et al. (2024a) proposed GaLore, a memory-efficient optimization algorithm that uses a low-rank projection of gradient matrices ${\bm{G}}$. Every $T$ steps, the current gradient matrix ${\bm{G}}_{t}$ is used to compute the projection matrix ${\bm{P}}$ via SVD decomposition. The gradient is then projected onto the low-rank space, where the optimization step is performed. Subsequently, the resulting low-rank update is projected back into the full-rank space and added to the weights ${\bm{W}}$.
  • BAdam. Luo et al. (2024) proposed a block coordinate descent (BCD)-type optimization method termed BAdam. The parameters are divided into blocks, which are then updated one by one using Adam. Similar to GaLore, the optimized block is updated every $T$ steps. Although this method was initially proposed only for fine-tuning, it is the closest method to our FRUGAL. Unlike BAdam, in our algorithm, state-free blocks are not frozen but are updated using signSGD.
  • Other Algorithms. Among other relevant methods, ReLoRA (Lialin et al., 2023) and MicroAdam (Modoranu et al., 2024) can also be highlighted. However, we did not include them for comparison in this paper for the following reasons: 1. ReLoRA was evaluated in (Zhao et al., 2024a), where it significantly underperformed compared to GaLore with the same memory budget. 2. MicroAdam. Its current implementation only supports bfloat16 master weights, whereas our main experiments conducted with mixed precision.
Main results. The results of our experiments are presented in Table 2, which includes both validation perplexity and memory footprint estimations for each method. We compared all memory-efficient methods under the same memory budget, specifically with a density $\rho=0.25$. Here, $\rho$ refers to the proportion of state-full parameters (see Section A.1 for a detailed explanation).
We conducted a grid search to determine the optimal learning rate for Adam, which we then applied uniformly to FRUGAL and BAdam (Luo et al., 2024). For GaLore (Zhao et al., 2024a), we found that using this same learning rate produced better results than the rate originally suggested in their paper. This discrepancy might be attributed to our experiments involving a significantly larger number of training steps than those for which GaLore’s original learning rate was optimized.
Table 2 demonstrates that FRUGAL significantly outperforms the memory-efficient baselines across all model sizes with the same memory budget, coming close to the performance of Adam.
Table 3: Perplexity of LLaMA-130M models pre-trained on C4 for 100k iterations (10B tokens). The leftmost column indicates the modules moved to the state-free set and trained using signSGD. The results show that Logits, unlike Embeddings and RMSNorms, are exceptionally responsive to the choice of optimization algorithm from Adam to signSGD.
State-free modules Perplexity $\downarrow$
Linear (corresponds to the FRUGAL with $\rho=0.0$) 20.02
Linear, RMSNorms 20.07
Linear, Embeddings 20.48
Linear, Embeddings, RMSNorms 20.55
Linear, Logits 34.66
Zero-density training. Table 2 also reveals a surprising result: FRUGAL with $\rho=0.0$ significantly outperforms both GaLore and BAdam, even when these competing methods use a higher density of $\rho=0.25$. Essentially, in this case, the parameters are divided into two parts — a state-full part consisting of the Embeddings, RMSNorms, and Logits, and a state-free part consisting of all other parameters. This division remains fixed throughout the training. We conducted additional experiments to determine the maximum subset of parameters that can be trained with a state-free optimizer without significant quality degradation. We systematically moved different combinations of the Embeddings, RMSNorms, and Logits from the state-full to the state-free set and observed the results during the training of LLaMA-130M. Table 3 reveals that the Logits demonstrates a dramatically higher sensitivity, with changes to its optimizer resulting in severe performance degradation. This finding aligns with results from (Zhao et al., 2024b), where the authors demonstrated that most parameters can be trained using SGDM, but the Logits require training with Adam.
5.1.2 Ablation study
We also conducted additional experiments to verify the robustness of our framework to various hyperparameters. Firstly, an ablation study on the state-full subspace update frequency $T$ in Table 10 shows that the performance keeps improving up to $T=200$. We note that, unlike in Zhao et al. (2024a), the perplexity does not significantly decrease even when reducing the update frequency to $T=10$ ($\sim 0.2$ drop vs. $\sim 4.$ drop for GaLore). A detailed explanation for this result can be found in Appendix C. Second, when using other schedulers, the performance gap between FRUGAL and baselines remains consistent, as shown in Tables 6 and 6. Then, the results of training in pure bfloat16 are presented in Table 8, demonstrating consistency with our main experiments in Table 2, i.e., FRUGAL significantly outperforms the baselines across these variations. We also conducted experiments to show how perplexity changes with varying $\rho$, and the results are presented in Table 11. Finally, we conducted an experiment to compare different strategies for selecting state-full blocks during training. The results in Table 10 show that there is no significant difference between random and structured block selection.
5.2 Fine-tuning experiments
Table 4: Evaluating FRUGAL for memory-efficient fine-tuning RoBERTa-Base on GLUE benchmark. Results represent the mean and standard deviation across 3 independent runs. Upper $\uparrow$ is better.
Method Modules Rank CoLA STS-B MRPC RTE SST2 MNLI QNLI QQP Avg
LoRA QV 8 63.8$\pm$.6 90.9$\pm$.1 89.1$\pm$.4 79.2$\pm$1.1 94.8$\pm$.2 87.6$\pm$.2 93.1$\pm$.1 90.6$\pm$.0 86.1
GaLore All 8 60.0$\pm$.2 90.8$\pm$.1 89.0$\pm$.7 79.7$\pm$.9 94.9$\pm$.5 87.6$\pm$.1 93.3$\pm$.1 91.1$\pm$.1 85.8
GaLore QV 8 56.1$\pm$.8 90.8$\pm$.2 88.1$\pm$.3 74.7$\pm$1.9 94.3$\pm$.1 86.6$\pm$.1 92.6$\pm$.1 89.4$\pm$.1 84.1
FRUGAL QV 8 64.5$\pm$.7 91.1$\pm$.1 89.2$\pm$.3 82.4$\pm$.9 94.8$\pm$.2 87.4$\pm$.1 92.8$\pm$.1 91.4$\pm$.1 86.7
FRUGAL None 0 64.8$\pm$.5 91.1$\pm$.1 89.1$\pm$.3 81.6$\pm$.6 94.9$\pm$.2 87.3$\pm$.1 92.8$\pm$.1 91.3$\pm$.1 86.6
We evaluated the performance of our framework in the context of memory-efficient fine-tuning using the GLUE benchmark (Wang, 2018), a widely-used collection of tasks for evaluating language models. Following the approach from Zhao et al. (2024a), we fine-tuned RoBERTa-base (Liu, 2019) using LoRA (Hu et al., 2021) and GaLore as baselines for comparison. We adhered to the setup described in LoRA, where low-rank updates of rank 8 were applied only to the Q and V matrices. For a detailed description of the experimental setup, see Section A.2.
However, this comparison required a minor modification to FRUGAL compared to the pre-training phase. Instead of selecting active parameters blockwise, we opted for columnwise selection in each matrix. This adjustment was necessary to ensure a fair comparison within a similar memory budget, as the number of trainable parameters in LoRA with rank 8 is approximately $2.5$ times fewer than the number of trainable parameters in any RoBERTa matrix. This transition from blockwise to columnwise selection allowed us to maintain comparable memory usage across methods. For the same reason, we did not include comparisons with BAdam (Luo et al., 2024) in this setup.
The results are presented in Table 4. Since the LoRA setup adds trainable adapters only to the Q and V matrices, while the GaLore code uses all modules as projectable parameters, we conducted experiments in both setups. The results demonstrate that FRUGAL significantly outperforms GaLore and shows comparable results to LoRA.
As in Section 5.1.1, we conducted additional experiments with FRUGAL using $\rho=0.0$. In this setup, only the classification head is trained using Adam, while the embedding parameters remain frozen, and the remaining parameters are trained using signSGD. The results demonstrate that this training approach barely compromises performance compared to FRUGAL with rank 8, and still outperforms GaLore. Similar to our findings in Section 5.1.1, we observe that the classification head parameters are particularly sensitive to the choice of optimizer, which can be seen in Table 13 where the model’s performance significantly deteriorates when using signSGD for classification head optimization.
6 Conclusion
In this work, we introduce a new memory-efficient optimization framework, FRUGAL. Within this framework, the optimization space is divided into two subspaces: the first is updated using a state-full algorithm such as Adam, while the second is updated using a state-free algorithm such as signSGD. We prove theoretical convergence guarantees for our framework with SGDM serving as the state-full algorithm and SGD as the state-free algorithm. In experiments involving pre-training and fine-tuning of language models, FRUGAL outperforms other approaches while using the same or smaller memory.
7 Acknowledgments
The work was done in the Laboratory of Federated Learning Problems of the ISP RAS (Supported by Grant App. No. 2 to Agreement No. 075-03-2024-214). The work was partially conducted while Philip Zmushko was a visiting research assistant in Mohamed bin Zayed University of Artificial Intelligence (MBZUAI).
Appendix A Additional experimental details
This section describes the main setups used in the experiments and presents additional experiments.
To begin, we introduce the hyperparameter density $\rho$. This hyperparameter represents the fraction of the total space in Linear layers that is updated with a stateful optimizer. For GaLore, this parameter is equal to $\rho=r/h$, where $r$ is the projection rank, and $h$ is the hidden size of the model. For RandK projection, this parameter can be expressed as $1-s$, where $s$ means sparsity. For BAdam and FRUGAL with the blockwise update, this parameter denotes the ratio of the number of active blocks $a_{\text{block}}$ to the total number of blocks $p$, i.e., $\rho=a_{\text{block}}/p$. When using FRUGAL with the column-wise update, as in Section 5.2, $\rho$ is equal to the ratio of the number of active columns $a_{\text{column}}$ to their total number $h$, i.e., $\rho=a_{\text{column}}/h$.
A.1 Pre-training
Setup. We adopt a LLaMA-based architecture with RMSNorm and SwiGLU (Wang, 2018) activations on the C4 dataset. Following Zhao et al. (2024a), we trained using a batch size of 512 sequences, sequence length of 256, weight decay of 0, and no gradient clipping. We used T5 tokenizer, since it also was trained on C4 with dictionary size equal to 32k. Additionally, the update frequency $T$ is set to 200, and the density $\rho=r/h$ is $0.25$ unless stated otherwise. We also should point out that similarly to Zhao et al. (2024a), we keep Embeddings, RMSNorms (Zhang & Sennrich, 2019), and Logits in the state-full subspace throughout the training and don’t reset the optimizer state for them.
We used standard Adam hyperparameters: $\beta_{1}=0.9,\beta_{2}=0.999,\varepsilon=1e-8$. For all the methods except GaLore, we selected the learning rate equal to the optimal learning rate for Adam, which we determined through a grid search among values $[1e-4,3e-4,1e-3,3e-3]$. FRUGAL’s learning rate for the state-free optimizer was set equal to that for the state-full optimizer for simplicity and ease of tuning. For a fair comparison with GaLore (Zhao et al., 2024a), we conducted experiments with two learning rate values: 1) the one specified by the authors in the original paper, and 2) the optimal learning rate for Adam, as used for other methods. We did this because the learning rate in the original paper could have been optimized for a different number of iterations.
To match the learning rate changes in the first steps of our training with Zhao et al. (2024a), we used a cosine learning rate schedule with restarts, with a warmup of 10% of the steps in a cycle length, and decay of the final learning rate down to 10% of the peak learning rate. To verify that our results are not sensitive to the choice of scheduler, we repeated the experiments for LLaMA-130M with other schedulers. Results for constant with warm-up and cosine (one cycle) with warm-up schedulers can be found in Tables 6 and 6.
Table 5: Perplexity of LLaMA-130M models pre-trained on C4 using constant scheduler with warm-up at various training iterations.
Method 100k 200k
Adam 19.51 18.51
GaLore, $\rho=0.25$ 22.63 21.03
BAdam, $\rho=0.25$ 22.31 20.66
FRUGAL, $\rho=0.25$ 19.97 18.85
FRUGAL, $\rho=0.0$ 20.33 19.14
Table 6: Perplexity of LLaMA-130M models pre-trained on C4 using cosine scheduler with warm-up at various training iterations.
Method 100k 200k
Adam 19.38 17.95
GaLore, $\rho=0.25$ 22.30 20.60
BAdam, $\rho=0.25$ 22.35 20.07
FRUGAL, $\rho=0.25$ 19.62 18.16
FRUGAL, $\rho=0.0$ 19.83 18.34
Table 7: Perplexity of LLaMA-130M models pre-trained on C4 using pure bfloat16 format both for model weights and optimizer statistics.
Method 100k iterations
Adam 21.88
GaLore, $\rho=0.25$ 24.19
BAdam, $\rho=0.25$ 25.03
FRUGAL, $\rho=0.25$ 23.17
FRUGAL, $\rho=0.0$ 22.64
Table 8: Perplexity of LLaMA-130M models pre-trained on C4 for 20k iterations (2.1B tokens) using SGD and signSGD with different learning rates. $\infty$ means that run diverged. LR stands for learning rate.
SGD signSGD
LR Perplexity LR Perplexity
0.1 184.83 3e-4 40.22
0.3 91.23 1e-3 41.18
1.0 $\infty$ 3e-3 109.32
Table 9: Perplexity of LLaMA-130M models pre-trained on C4 for 200k iterations using FRUGAL with $\rho=1/3$ and different Block update strategy, taken from Luo et al. (2024).
Method Perplexity
Random 18.50
Ascending 18.54
Descending 18.50
Table 10: Perplexity of LLaMA-130M models pre-trained on C4 for 200k iterations (20B tokens) using FRUGAL with $\rho=0.25$ and different update frequency $T$.
Update frequency $T$ Perplexity
10 18.82
20 18.73
50 18.69
100 18.65
200 18.60
500 18.60
1000 18.61
Table 11: Perplexity of LLaMA-130M models pre-trained on C4 for 200k iterations (20B tokens) using FRUGAL with different density $\rho$.
FRUGAL
$\rho$ 1.0 (Adam) 0.5 0.33 0.25 0.125 0.0625 0.0 signSgd
Perplexity 18.13 18.40 18.50 18.63 18.71 18.80 18.90 33.22
A.2 Fine-tuning
Setup. The batch size and learning rate values used for FRUGAL in the experiments from Table 4 are presented in Table 12. In all experiments, we set the learning rate for the state-free optimizer to $1/10$ of the learning rate of the state-full optimizer. Other hyperparameters, such as scheduler, number of epochs, maximum sequence length, and warmup ratio, were taken from Hu et al. (2021).
Table 12: Hyperparameters of fine-tuning RoBERTa-base for FRUGAL.
MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
Batch Size 128 128 16 256 256 128 32 16
State-full Learning Rate 5E-05 5E-05 2E-04 5E-04 1E-04 5E-05 2E-04 1E-04
State-free lr multiplier 0.1
Rank/Density $r=8\;\;/\;\;r=0\;(\rho=0)$
We also present a comparison between fine-tuning using FRUGAL with $\rho=0.0$ and full fine-tuning using signSGD. Essentially, the only difference is that in the second case, the classification head is updated with signSGD instead of Adam. The results in Table 13 show that the classification head is extremely sensitive to the optimizer type, and switching the optimizer significantly drops the accuracy.
Table 13: Results of fine-tuning RoBERTa-Base on several tasks from GLUE. The left column indicates which modules were trained using the state-full optimizer Adam. The remaining modules, except for the frozen Embedding layer, were trained using the state-free signSGD.
Method SST2 QNLI QQP
Classification head (corresponds to the FRUGAL with $\rho=0.0$) 94.9$\pm$.2 92.8$\pm$.1 91.3$\pm$.1
None (corresponds to the fine-tuning using signSGD) 89.7 81.6 74.3
Appendix B Memory estimation
In this section, we will examine memory requirements for different projection types using the LLaMA-like architecture as an example and show that RandK, column-wise, and blockwise projections result in approximately the same amount of additional memory for a given density value $\rho$ Appendix A. In contrast, the semi-orthogonal projection matrix (GaLore-like) requires a slightly larger value in this setup. Recall that we follow the setup from Zhao et al. (2024a), where Embeddings, RMSNorms, and Logits remain in the state-full subspace throughout the training, so the projection does not interact with them, and they give the same memory overhead for all projection methods.
Let the number of parameters in the remaining projectable parameters be $P$. Then, training using Adam gives an additional overhead of $2P$ float values for storing $m$ and $v$ for each parameter. Now, let’s consider blockwise and column-wise projections and suppose we want to achieve a density $\rho$. For blockwise, we take $\text{round}(\rho\cdot L)$ layers, where $L$ is the total number of transformer layers, and for column-wise, we take $\text{round}(\rho\cdot k)$ columns for each matrix of size $n\times k$. Since the memory required to store block or column indices is negligible compared to other costs, we find that the total size of the optimizer state when using Adam as a state-full optimizer will be $2\rho\cdot P$, with an adjustment for rounding.
In the case of RandK projection, we have the same $2\rho\cdot P$ float values ${\bm{M}}$ and ${\bm{V}}$ in the optimizer state. However, we must also know the current indices corresponding to these values. On the other hand, it is widely known that if one needs to save a set of random values, they don’t need to store all these values - it’s sufficient to store only the seed from which they were generated. Thus, for RandK, the total memory also equals $2\rho\cdot P$.
If we recalculate this considering a specific LLaMA-like architecture, each layer consists of 7 matrices: 4 matrices of size $h\times h$ (Query, Key, Value, Output) and 3 matrices of size $h\times h_{ff}$ (Gate, Down, Up), where $h$ is the hidden size of the model, and $h_{ff}$ is the FFN hidden size. In the LLaMA architecture, it’s typically:
$h_{ff}=4h\cdot\frac{2}{3}=\frac{8}{3}h.$
Then, the amount of memory for RandK projection (and consequently for all others mentioned above) is:
$2\cdot(4\cdot(\rho h^{2})+3\cdot(\rho\cdot h\cdot h_{ff}))=2\cdot(4\cdot\rho h^{2}+3\cdot(\frac{8}{3}\rho\cdot h^{2}))=24\rho\cdot h^{2}$
for each layer on average ($2$ corresponds to the number of matrices ${\bm{M}}$ and ${\bm{V}}$).
In the case of a GaLore-like semi-orthogonal projection matrix, the situation is as follows. We have projections onto a low-rank subspace of rank $r$, where $r=\text{round}(\rho\cdot h)$. Then, for Query, Key, Value, and Output projections, we need to store ${\bm{P}},{\bm{M}},{\bm{V}}\in\mathbb{R}^{h\times r},$ and for Gate, Down and Up projections either ${\bm{P}}\in\mathbb{R}^{h\times r},{\bm{M}},{\bm{V}}\in\mathbb{R}^{h_{ff}\times r}$, or ${\bm{P}}\in\mathbb{R}^{h_{ff}\times r},{\bm{M}},{\bm{V}}\in\mathbb{R}^{h\times r}$. Since the second option requires less memory, it is used by default in (Zhao et al., 2024a) and, therefore, in FRUGAL, too. Then, the total memory requirements are:
$4\cdot(3\cdot rh)+3\cdot(2\cdot r\cdot h+r\cdot h_{ff})=12rh+6rh+3rh_{ff}=(12+6+3\cdot\frac{8}{3})rh=26\rho h^{2}.$
To sum up, RandK, column-wise and blockwise projection requires $2\rho P$ additional memory, while semi-orthogonal projection (GaLore-like) requires $\frac{26}{24}\cdot 2\rho P=\frac{13}{12}\cdot 2\rho P$ additional memory.
Let’s recall that in addition to this, SVD requires additional computation, which can take up to 10% as the model size increases (Zhao et al., 2024a). Therefore, for our method, we settled on blockwise projection.
Appendix C Optimizer state management
In this section, we would like to propose some modifications to the GaLore algorithm. These modifications are also used in our framework as SVD projection.
Specifically, we want to consider the projection of the state when changing the active subspace. In GaLore (Zhao et al., 2024a), when updating the projection, the optimizer states $M$ and $V$ do not change. This results in new projected gradients and old $M$ and $V$ being in different subspaces. This implementation has little effect on the result with large values of update frequency $T$, as the values of $M$ and $V$ from the previous subspace decay exponentially quickly. However, more frequent changes $T$ significantly affect the result. We hypothesize that this is why in Zhao et al. (2024a) the model quality degraded so significantly when $T$ was decreased, while as seen in Table 10, FRUGAL experiences much less degradation.
There are two different ways to overcome this obstacle: either project the state back to full-rank space or reset the state before a new round. However, the first option may be challenging in the case of arbitrary projection. Specifically, while it’s possible to project momentum back to full-rank space (see Alg. 2 in Hao et al. (2024)), the same cannot be easily done with variance because its values depend quadratically on the projection matrix. However, the projection of variance will also be trivial if the set of basis vectors for the projection is fixed, which is true, for example, for coordinate projection with RandK.
Figure 3: Toy example of solving quadratic minimization problem with GaLore-like SGDM with and without re-projection of optimizer state. Algorithm with re-projection converges much faster.
To demonstrate the effectiveness of this improvement, we provide a toy example. We consider a quadratic minimization problem of $\|W\|^{2},W\in\mathbb{R}^{10\times 10}$. For optimization, we use GaLore-like SGDM and GaLore-like SGDM with Momentum state projection. This projection is similar to Alg. 2 from (Hao et al., 2024), except we additionally normalize the new momentum by the ratio of norms before and after re-projection to preserve momentum mass. We use ranks of $3$ and $6$, and an update frequency $T=10$ and plot mean and standard deviation across $5$ independent runs. The results are presented in Figure 3. As can be seen, the variant with state projection converges much faster.
Appendix D LLaMA 7B pre-training results.
In this section, we present the results of pre-training a LLaMA 7b model on the C4 dataset for 120k iterations on 12B tokens. See results in Table 14. We conducted the training in pure bfloat16 with the density $\rho=0.0$. We used learning rate $0.0005$ for state-full optimizer and $0.00015$ for state-free optimizer. However, unlike Zhao et al. (2024a), we didn’t use Adam8bit for state-full parameters but rather Adam, so it may not be an entirely fair comparison. Nevertheless, the results show that FRUGAL has the potential for scaling up to 7B parameter models.
Table 14: Pre-training LLaMA 7B on C4 dataset for 120K steps. Validation perplexity and memory estimate are reported. We also indicate the additional memory overhead introduced by the optimization algorithm.
40K 80K 120K
8-bit Adam 18.09 15.47 14.83
8-bit GaLore 17.94 15.39 14.95
FRUGAL, $\rho=0.0$ 17.56 14.50 13.49
Tokens (B) 5.2 10.5 15.7
Appendix E Convergence Theory
Firstly, we provide ommited definition of $L$-smooth function.
Definition 1.
We say that $f:\operatorname*{\mathbb{R}^{d}}\rightarrow\mathbb{R}$ is $L-$smooth with $L\geq 0$, if it is differentiable and satisfies
$f(y)\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{L}{2}\|y-x\|^{2},\forall x,y\in\operatorname*{\mathbb{R}^{d}}.$
Below, we provide an equivalent formulation of Algorithm 2 that enables us to use the proof of the similar structure to SGDM momentum analyis of Liu et al. (2020).
Algorithm 3 FRUGAL(SGDM, SGD): Equivalent to Algorithm 2 for constant step isze
Input: momentum weight $\beta\in[0,1)$, initialization $x^{1}\in\operatorname*{\mathbb{R}^{d}}$ and $m^{0}=0$, step sizes $\{\alpha_{k}:=\alpha>0\}_{k=1}^{K}$, momentum set $J_{k}\subset[d]$ for $k=1,2\ldots$.
1:for $k=1,2,\ldots$ do
2:     Compute stochastic gradient $\tilde{g}^{k}\leftarrow\nabla f_{\zeta^{k}}(x^{k})$;
3:     Update momentum vector $\tilde{m}_{j}^{k}\leftarrow(1-\beta)\tilde{g}_{j}^{k}+\beta\begin{cases}\tilde{m}_{j}^{k-1}&\text{if }j\in J_{k},\\ 0&\text{otherwise;}\end{cases}$
4:     Update iterate $x^{k+\nicefrac{{1}}{{2}}}\leftarrow x^{k}-\alpha\tilde{m}^{k}$;
5:     $x_{j}^{k+1}\leftarrow\begin{cases}\frac{x_{j}^{k+\nicefrac{{1}}{{2}}}}{1-\beta}-\frac{\beta x^{k}_{j}}{1-\beta}&\text{if }j\notin J_{k+1},\\ x_{j}^{k+\nicefrac{{1}}{{2}}}&\text{otherwise;}\end{cases}$
6:end for
Next, we present several key ingredients of the proof. Firstly, we can express the momentum term $\tilde{m}_{j}^{k}$ as
$\tilde{m}_{j}^{k}$ $=(1-\beta)\sum_{i=t^{k}_{j}}^{k}\beta^{k-i}\tilde{g}_{j}^{i},$ (3)
where $t^{k}_{j}:=\max_{t\leq k}\{j\notin J_{t}\}$, i.e., the last time when the momentum buffer was released. We denote
$m_{j}^{k}$ $=(1-\beta)\sum_{i=t^{k}_{j}}^{k}\beta^{k-i}g_{j}^{i},$ (4)
Using this notation, we proceed with two lemmas, one showing variance reduction effect of momentum, the other boundess of momentum bias.
Lemma 1.
Under Assumption 1, the update vector $\tilde{m}^{k}$ in Algorithm 3 satisfies
$\mathbb{E}\left[\left\|\tilde{m}^{k}-m^{k}\right\|^{2}\right]$ $\leq\frac{1-\beta}{1+\beta}\sigma^{2}.$
Proof.
Since $\tilde{m}_{j}^{k}=(1-\beta)\sum_{i=t_{j}^{k}}^{k}\beta^{k-i}\tilde{g}_{j}^{i}$, we have
$\mathbb{E}\left[\left\|\tilde{m}^{k}-m^{k}\right\|^{2}\right]$ $=\sum_{j\in[d]}\mathbb{E}\left[\left\|\tilde{m}_{j}^{k}-m_{j}^{k}\right\|^{2}\right]$
$\leq(1-\beta)^{2}\sum_{j\in[d]}\mathbb{E}\left[\left\|\sum_{i=t^{k}_{j}}^{k}\beta^{k-i}(\tilde{g}_{j}^{i}-g_{j}^{i})\right\|^{2}\right].$
Moreover, since $\zeta^{1},\zeta^{2},...,\zeta^{k}$ are independent random variables (item 3 of Assumption 1), we can use conditional expectation to show that $\mathbb{E}\left[(\tilde{g}_{j}^{i_{1}}-g_{j}^{i_{1}})(\tilde{g}_{j}^{i_{2}}-g_{j}^{i_{2}})\right]=0$ for $i_{1}\neq i_{2}$. Therefore,
$\mathbb{E}\left[\left\|\tilde{m}^{k}-m^{k}\right\|^{2}\right]$ $\leq(1-\beta)^{2}\sum_{j\in[d]}\mathbb{E}\left[\sum_{i=t^{k}_{j}}^{k}\beta^{2(k-i)}\|\tilde{g}^{i}_{j}-g^{i}_{j}\|^{2}\right]$
$\leq\frac{1-\beta}{1+\beta}\sum_{j\in[d]}\mathbb{E}\left[(1-\beta^{2(k-t^{k}_{j}+1)})\right]\sigma_{j}^{2}$
$\leq\frac{1-\beta}{1+\beta}\sum_{j\in[d]}\sigma_{j}^{2}=\frac{1-\beta}{1+\beta}\sigma^{2}.$
Lemma 2.
Under Assumption 1, the update vector $\tilde{m}^{k}$ in Algorithm 3 further satisfies
$\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]\leq p^{k}_{\max}\mathbb{E}\left[\sum_{i=1}^{k-1}a_{k,i}\|x^{i+1}-x^{i}\|^{2}\right],$
where $k_{j}=k-t^{k}_{j}+1$, and
$a_{k,i}=L^{2}\beta^{k-i}\left(k-i+\frac{\beta}{1-\beta}\right).$ (5)
Proof.
Let $\Pr_{k-1}[j\in J_{k}]=p^{k}_{j}$ and $p^{k}_{\max}:=\max_{j\in[d]}\{p^{k}_{j}\}$. Then,
$\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]$ $=\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{1-\beta}{1-\beta^{k_{j}}}\sum_{i=t^{k}_{j}}^{k}\beta^{k-i}(g^{i}_{j}-g_{j}^{k})\right\|^{2}\right]$
$=(1-\beta)^{2}\mathbb{E}\left[\sum_{j\in J_{k}}\sum_{i,l=t^{k}_{j}}^{k}\langle\beta^{k-i}(g^{k}_{j}-g^{i}_{j}),\beta^{k-l}(g^{k}_{j}-g^{l}_{j})\rangle\right]$
$\leq(1-\beta)^{2}\mathbb{E}\left[\sum_{j\in J_{k}}\sum_{i,l=1}^{k}\beta^{2k-i-l}\left(\frac{1}{2}\|g_{j}^{k}-g_{j}^{i}\|^{2}]+\frac{1}{2}\|g_{j}^{k}-g_{j}^{l}\|^{2}\right)\right]$
$=(1-\beta)^{2}\mathbb{E}\left[\sum_{j\in J_{k}}\sum_{i=1}^{k}\left(\sum_{l=1}^{k}\beta^{2k-i-l}\right)\frac{1}{2}\mathbb{E}[\|g^{k}_{j}-g^{l}_{j}\|^{2}\right]$
$\quad+(1-\beta)^{2}\mathbb{E}\left[\sum_{j\in J_{k}}\sum_{l=1}^{k}\left(\sum_{i=1}^{k}\beta^{2k-i-l}\right)\frac{1}{2}[\|g_{j}^{k}-g_{j}^{i}\|^{2}\right]$
$=(1-\beta)^{2}\mathbb{E}\left[\sum_{j\in J_{k}}\sum_{i=1}^{k}\frac{\beta^{k-i}(1-\beta^{k_{j}})}{1-\beta}\|g_{j}^{k}-g_{j}^{i}\|^{2}\right]$
$\leq(1-\beta)\mathbb{E}\left[\sum_{j\in J_{k}}\sum_{i=1}^{k}\beta^{k-i}\|g_{j}^{k}-g_{j}^{i}\|^{2}\right],$
$\leq(1-\beta)p^{k}_{\max}\mathbb{E}\left[\sum_{i=1}^{k}\beta^{k-i}\|g^{k}-g^{i}\|^{2}\right],$
where we applied Cauchy-Schwarz to the first inequality.
By applying triangle inequality and the smoothness of $f$ (item 1 in Assumption 1), we further have
$\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]$ $\leq(1-\beta)p^{k}_{\max}\mathbb{E}\left[\sum_{i=1}^{k}\beta^{k-i}(k-i)\sum_{l=i}^{k-1}\|g^{l+1}-g^{l}\|^{2}\right]$
$\leq\mathbb{E}\left[\sum_{l=1}^{k-1}\left((1-\beta)p^{k}_{\max}L^{2}\sum_{i=1}^{l}\beta^{k-i}(k-i)\right)\|x^{l+1}-x^{l}\|^{2}\right].$
Therefore, by defining $a^{\prime}_{k,l}=(1-\beta)L^{2}\sum_{i=1}^{l}\beta^{k-i}(k-i)$, we get
$\begin{split}&\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]\leq p^{k}_{\max}\mathbb{E}\left[\sum_{l=1}^{k-1}a^{\prime}_{k,l}\|x^{l+1}-x^{l}\|^{2}\right].\end{split}$ (6)
Furthermore, $a^{\prime}_{k,j}$ can be calculated as
$\begin{split}a^{\prime}_{k,l}&=L^{2}\beta^{k}\left(-(k-1)-\frac{1}{1-\beta}\right)+L^{2}\beta^{k-l}\left(k-l+\frac{\beta}{1-\beta}\right).\end{split}$ (7)
Notice that
$a^{\prime}_{k,l}<a_{k,l}\coloneqq L^{2}\beta^{k-l}\left(k-l+\frac{\beta}{1-\beta}\right).$ (8)
Combining this with equation 6, we arrive at
$\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]\leq p^{k}_{\max}\mathbb{E}\left[\sum_{i=1}^{k-1}a_{k,i}\|x^{i+1}-x^{i}\|^{2}\right],$
where
$a_{k,i}=L^{2}\beta^{k-i}\left(k-i+\frac{\beta}{1-\beta}\right).$
From Lemma 2, we know that the distance of the non-stochastic momentum from $g^{k}$ is bounded by the weighted sum of past successive iterate differences. Furthermore, the coefficients $a_{k,i}$ decays exponentially in $\beta$.
Therefore, we use the following Lyapunov function
$L^{k}=\left(f(z^{k})-f^{\star}\right)+\sum_{i=1}^{k-1}c_{i}\|x^{k+1-i}-x^{k-i}\|^{2}.$ (9)
for some positive $c_{i}$ that we specify later. As it is common for convergence theory of SGDM to analyze an auxiliary sequence $z^{k}$ defined as
$z_{j}^{k}=\begin{cases}x_{j}^{k}&k=1,\\ \frac{1}{1-\beta}x_{j}^{k-\nicefrac{{1}}{{2}}}-\frac{\beta}{1-\beta}x_{j}^{k-1}&\,k\geq 2,\end{cases}$ (10)
which behaves more like an SGD iterate, although the stochastic gradient $\tilde{g}^{k}$ is not taken at $z^{k}$.
Lemma 3.
Let $x^{k}$’s be iterates of Algorithm 3, then $z^{k}$ defined in equation 10 satisfies
$z^{k+1}-z^{k}$ $=-\alpha\tilde{g}^{k}.$
Proof.
We have to consider two different cases. Firstly, if $k=1$ or $j\notin J_{k}$, then
$z^{k+1}_{j}-z^{k}_{j}$ $=\frac{x_{j}^{k+\nicefrac{{1}}{{2}}}}{1-\beta}-\frac{\beta x^{k}_{j}}{1-\beta}-x^{k}_{j}=\frac{x^{k}_{j}-\alpha\tilde{m}_{j}^{k}-\beta x^{k}_{j}-(1-\beta)x^{k}_{j}}{1-\beta}=-\frac{\alpha(1-\beta)\tilde{g}_{j}^{k}}{1-\beta}=-\alpha\tilde{g}^{k}_{j}.$
Secondly, if $k\geq 2$, $j\in J_{k}$, then
$z_{j}^{k+1}-z_{j}^{k}$ $=\frac{1}{1-\beta}(x_{j}^{k+\nicefrac{{1}}{{2}}}-x_{j}^{k-\nicefrac{{1}}{{2}}})-\frac{\beta}{1-\beta}(x_{j}^{k}-x_{j}^{k-1})$
$=\frac{1}{1-\beta}(x_{j}^{k+\nicefrac{{1}}{{2}}}-x_{j}^{k})-\frac{\beta}{1-\beta}(x_{j}^{k}-x_{j}^{k-1})$
$=\frac{1}{1-\beta}(-\alpha\tilde{m}_{j}^{k})-\frac{\beta}{1-\beta}(-\alpha\tilde{m}_{j}^{k-1})$
$=\frac{1}{1-\beta}(-\alpha\tilde{m}_{j}^{k}+\alpha\beta\tilde{m}_{j}^{k-1})=-\alpha\tilde{g}_{j}^{k}.$
Before procceding with the main convergence theory, we require one more proposition that shows descent in objective value.
Proposition 1.
Take Assumption 1. Then, for $z^{k}$ defined in equation 10, we have
$\begin{split}\mathbb{E}[f(z^{k+1})]&\leq\mathbb{E}[f(z^{k})]+\left(-\alpha+\frac{1+\beta^{2}}{1-\beta}L\alpha^{2}+\frac{1}{2}L\alpha^{2}\right)\mathbb{E}[\|g^{k}\|^{2}]\\ &\quad+\left(\frac{\beta^{2}}{2(1+\beta)}+\frac{1}{2}\right)L\alpha^{2}\sigma^{2}+\frac{L\alpha^{2}}{1-\beta}\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right].\end{split}$ (11)
Proof.
The smoothness of $f$ yields
$\begin{split}\mathbb{E}_{\zeta^{k}}[f(z^{k+1})]&\leq f(z^{k})+\mathbb{E}_{\zeta^{k}}[\langle\nabla f(z^{k}),z^{k+1}-z^{k}\rangle]+\frac{L}{2}\mathbb{E}_{\zeta^{k}}[\|z^{k+1}-z^{k}\|^{2}]\\ &=f(z^{k})+\mathbb{E}_{\zeta^{k}}[\langle\nabla f(z^{k}),-\alpha\tilde{g}^{k}\rangle]+\frac{L\alpha^{2}}{2}\mathbb{E}_{\zeta^{k}}[\|\tilde{g}^{k}\|^{2}],\end{split}$ (12)
where we have applied Lemma 3 in the second step.
For the inner product term, we can take full expectation $\mathbb{E}=\mathbb{E}_{\zeta^{1}}...\mathbb{E}_{\zeta^{k}}$ to get
$\mathbb{E}[\langle\nabla f(z^{k}),-\alpha\tilde{g}^{k}\rangle]=\mathbb{E}[\langle\nabla f(z^{k}),-\alpha g^{k}\rangle],$
which follows from the fact that $z^{k}$ is determined by the previous $k-1$ random samples $\zeta^{1},\zeta^{2},...\zeta^{k-1}$, which is independent of $\zeta^{k}$, and $\mathbb{E}_{\zeta^{k}}[\tilde{g}^{k}]=g^{k}$.
So, we can bound
$\mathbb{E}[\langle\nabla f(z^{k}),-\alpha\tilde{g}^{k}\rangle]$ $=\mathbb{E}[\langle\nabla f(z^{k})-g^{k},-\alpha g^{k}\rangle]-\alpha\mathbb{E}[\|g^{k}\|^{2}]$
$\leq\alpha\frac{\rho_{0}}{2}L^{2}\mathbb{E}[\|z^{k}-x^{k}\|^{2}]+\alpha\frac{1}{2\rho_{0}}\mathbb{E}[\|g^{k}\|^{2}]-\alpha\mathbb{E}[\|g^{k}\|^{2}],$
where $\rho_{0}>0$ can be any positive constant (to be determined later).
Combining equation 12 and the last inequality, we arrive at
$\mathbb{E}[f(z^{k+1})]$ $\leq\mathbb{E}[f(z^{k})]+\alpha\frac{\rho_{0}}{2}L^{2}\mathbb{E}[\|z^{k}-x^{k}\|^{2}]$
$\quad+(\alpha\frac{1}{2\rho_{0}}-\alpha)\mathbb{E}[\|g^{k}\|^{2}]+\frac{L\alpha^{2}}{2}\mathbb{E}[\|\tilde{g}^{k}\|^{2}].$
By construction, $z_{j}^{k}-x_{j}^{k}=-\frac{\beta}{1-\beta}\alpha\tilde{m}_{j}^{k-1}$ for $j\in J_{k}$, 0 otherwise. Consequently,
$\begin{split}\mathbb{E}[f(z^{k+1})]&\leq\mathbb{E}[f(z^{k})]+\alpha^{3}\frac{\rho_{0}}{2}L^{2}(\frac{\beta}{1-\beta})^{2}\mathbb{E}\left[\sum_{j\in J_{k}}\|\tilde{m}_{j}^{k-1}\|^{2}\right]\\ &\quad+(\alpha\frac{1}{2\rho_{0}}-\alpha)\mathbb{E}[\|g^{k}\|^{2}]+\frac{L\alpha^{2}}{2}\mathbb{E}[\|\tilde{g}^{k}\|^{2}].\end{split}$ (13)
Let $k_{j}=k-t^{k-1}_{j}+1$. Then, from Lemma 1 we know that
$\begin{split}\mathbb{E}\left[\sum_{j\in J_{k}}\|\tilde{m}_{j}^{k-1}\|^{2}\right]&\leq 2\mathbb{E}\left[\sum_{j\in J_{k}}\|\tilde{m}_{j}^{k-1}-m_{j}^{k-1}\|^{2}\right]+2\mathbb{E}\left[\sum_{j\in J_{k}}\|m_{j}^{k-1}\|^{2}\right]\\ &\leq 2\frac{1-\beta}{1+\beta}\mathbb{E}\left[\sum_{j\in J_{k}}\sigma_{j}^{2}+2\sum_{j\in J_{k}}\|m_{j}^{k-1}\|^{2}\right]\\ \mathbb{E}\left[\sum_{j\in J_{k}}\|m_{j}^{k-1}\|^{2}\right]&=\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{(k-1)_{j}})^{2}\left\|\frac{m_{j}^{k-1}}{(1-\beta^{(k-1)_{j}})}\right\|^{2}\right]\\ &\leq 2\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{(k-1)_{j}})^{2}\left\|\frac{m_{j}^{k-1}}{(1-\beta^{(k-1)_{j}})}-g_{j}^{k}\right\|^{2}\right]+2\mathbb{E}\left[\sum_{j\in J_{k}}\left\|g_{j}^{k}\right\|^{2}\right]\\ \mathbb{E}\left[\|\tilde{g}^{k}\|^{2}\right]&\leq\sigma^{2}+\mathbb{E}[\|g^{k}\|^{2}].\end{split}$ (14)
Putting these into equation 13, we arrive at
$\mathbb{E}[f(z^{k+1})]$ $\leq\mathbb{E}[f(z^{k})]+\bigg{(}-\alpha+\alpha\frac{1}{2\rho_{0}}+2\alpha^{3}\rho_{0}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{L\alpha^{2}}{2}\bigg{)}\mathbb{E}[\|g^{k}\|^{2}]$
$\quad+\left(\alpha^{3}{\rho_{0}}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}\frac{1-\beta}{1+\beta}\sigma^{2}+\frac{L\alpha^{2}}{2}\sigma^{2}\right)$
$\quad+2\alpha^{3}\rho_{0}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{(k-1)_{j}})^{2}\left\|\frac{m_{j}^{k-1}}{(1-\beta^{(k-1)_{j}})}-g_{j}^{k}\right\|^{2}\right].$
Notice that if $j\in J^{k}$, then $(k-1)_{j}=k_{j}-1$. Therefore,
$\mathbb{E}\left[\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]$ $=\mathbb{E}\left[\left\|\frac{\beta m_{j}^{k-1}+(1-\beta)g^{k}_{j}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]$
$=\beta^{2}\mathbb{E}\left[\left(\frac{(1-\beta^{k_{j}-1})}{(1-\beta^{k_{j}})}\right)^{2}\left\|\frac{m_{j}^{k-1}}{(1-\beta^{(k-1)_{j}})}-g_{j}^{k}\right\|^{2}\right].$
Substituting the above into the last inequality produces
$\begin{split}\mathbb{E}[f(z^{k+1})]&\leq\mathbb{E}[f(z^{k})]+\bigg{(}-\alpha+\alpha\frac{1}{2\rho_{0}}+2\alpha^{3}\rho_{0}L^{2}(\frac{\beta}{1-\beta})^{2}+\frac{L\alpha^{2}}{2}\bigg{)}\mathbb{E}[\|g^{k}\|^{2}]\\ &\quad+\left(\alpha^{3}{\rho_{0}}L^{2}(\frac{\beta}{1-\beta})^{2}\frac{1-\beta}{1+\beta}\sigma^{2}+\frac{L\alpha^{2}}{2}\sigma^{2}\right)\\ &\quad+2\alpha^{3}\rho_{0}L^{2}\left(\frac{1}{1-\beta}\right)^{2}\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right].\end{split}$ (15)
Finally, $\rho_{0}=\frac{1-\beta}{2L\alpha}$ gives
$\begin{split}\mathbb{E}[f(z^{k+1})]&\leq\mathbb{E}[f(z^{k})]+\left(-\alpha+\frac{1+\beta^{2}}{1-\beta}L\alpha^{2}+\frac{1}{2}L\alpha^{2}\right)\mathbb{E}[\|g^{k}\|^{2}]\\ &\quad+\left(\frac{\beta^{2}}{2(1+\beta)}+\frac{1}{2}\right)L\alpha^{2}\sigma^{2}+\frac{L\alpha^{2}}{1-\beta}\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right].\end{split}$
E.1 Convergence of Algorithm 3
Firstly, by combining results from prior section, we can bound our Lyapunov function $L^{k}$ defined in equation 9.
Proposition 2.
Let Assumption 1 hold and $\alpha\leq\frac{1-\beta}{2\sqrt{2}L\sqrt{p^{k}_{\max}}\sqrt{\beta+\beta^{2}}}$ in Algorithm 3. Let $\{c_{i}\}_{i=1}^{\infty}$ in equation 9 be defined by
$c_{1}$ $=\frac{\frac{\beta+\beta^{2}}{(1-\beta)^{3}}L^{3}\alpha^{2}}{1-4\alpha^{2}\frac{\beta+\beta^{2}}{(1-\beta)^{2}}L^{2}},\quad\quad c_{i+1}=c_{i}-\left(4c_{1}\alpha^{2}+\frac{L\alpha^{2}}{1-\beta}\right)\beta^{i}(i+\frac{\beta}{1-\beta})L^{2}\quad\text{for all $i\geq 1$.}$
Then, $c_{i}>0$ for all $i\geq 1$, and
$\begin{split}\mathbb{E}[L^{k+1}-L^{k}]&\leq\left(-\alpha+\frac{3-\beta+\beta^{2}}{2(1-\beta)}L\alpha^{2}+4c_{1}\alpha^{2}\right)\mathbb{E}[\|g^{k}\|^{2}]\\ &\quad+\left(\frac{\beta^{2}}{2(1+\beta)}L\alpha^{2}\sigma^{2}+\frac{1}{2}L\alpha^{2}\sigma^{2}+2c_{1}\alpha^{2}\sigma^{2}\right).\end{split}$ (16)
Proof.
Recall that $L^{k}$ is defined as
$L^{k}=f(z^{k})-f^{*}+\sum_{i=1}^{k-1}c_{i}\|x^{k+1-i}-x^{k-i}\|^{2},$
Therefore, by equation 15 we know that
$\begin{split}&\mathbb{E}[L^{k+1}-L^{k}]\leq\\ &\quad(-\alpha+\frac{1+\beta^{2}}{1-\beta}L\alpha^{2}+\frac{1}{2}L\alpha^{2})\mathbb{E}[\|g^{k}\|^{2}]\\ &\quad+\sum_{i=1}^{k-1}(c_{i+1}-c_{i})\mathbb{E}[\|x^{k+1-i}-x^{k-i}\|^{2}]+c_{1}\mathbb{E}[\|x^{k+1}-x^{k}\|^{2}]\\ &\quad+\left(\frac{\beta^{2}}{2(1+\beta)}+\frac{1}{2}\right)L\alpha^{2}\sigma^{2}+\frac{L\alpha^{2}}{1-\beta}\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right].\end{split}$ (17)
To bound the $c_{1}\mathbb{E}[\|x^{k+1}-x^{k}\|^{2}]$ term, we need the following inequalities, which are obtained similarly as equation 14.
$\begin{split}\mathbb{E}[\|\tilde{m}^{k}\|^{2}]&\leq 2\frac{1-\beta}{1+\beta}\sigma^{2}+2\mathbb{E}[\|m^{k}\|^{2}]\\ \mathbb{E}[\|m^{k}\|^{2}]&\leq 2\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]+2\mathbb{E}\left[\left\|g^{k}\right\|^{2}\right]\\ \mathbb{E}[\|\tilde{g}^{k}\|^{2}]&\leq\sigma^{2}+\mathbb{E}[\|g^{k}\|^{2}].\end{split}$ (18)
Let $\Pr_{k-1}[j\in J_{k}]=p^{k}_{j}$ and $p^{k}_{\min}:=\min_{j\in[d]}\{p^{k}_{j}\}$. Then, $c_{1}\mathbb{E}[\|x^{k+1}-x^{k}\|^{2}]$ can be bounded as
$c_{1}\mathbb{E}[\|x^{k+1}-x^{k}\|^{2}]$ $=c_{1}\alpha^{2}\mathbb{E}[\|\tilde{u}^{k}\|^{2}]=c_{1}\alpha^{2}\mathbb{E}\left[\sum_{j\in J_{k}}\|\tilde{m}_{j}^{k}\|^{2}+\sum_{j\notin J_{k}}\|\tilde{g}_{j}^{k}\|^{2}\right]$
$\leq c_{1}\alpha^{2}\mathbb{E}\left[\|\tilde{m}^{k}\|^{2}+(1-p^{k}_{\min})\|\tilde{g}^{k}\|^{2}\right]$
$\leq c_{1}\alpha^{2}\left(\left(2\frac{1-\beta}{1+\beta}+1-p^{k}_{\min}\right)\sigma^{2}+5\mathbb{E}[\|g^{k}\|^{2}]\right)$
$\quad+4c_{1}\alpha^{2}\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]$
Combine this with equation 17, we obtain
$\begin{split}&\mathbb{E}[L^{k+1}-L^{k}]\\ &\leq(-\alpha+\frac{1+\beta^{2}}{1-\beta}L\alpha^{2}+\frac{1}{2}L\alpha^{2}+5c_{1}\alpha^{2})\mathbb{E}[\|g^{k}\|^{2}]+\left(\frac{\beta^{2}}{2(1+\beta)}+\frac{1}{2}+\frac{c_{1}}{L}\left(2\frac{1-\beta}{1+\beta}+1-p^{k}_{\min}\right)\right)L\alpha^{2}\sigma^{2}\\ &\quad+\sum_{i=1}^{k-1}(c_{i+1}-c_{i})\mathbb{E}[\|x^{k+1-i}-x^{k-i}\|^{2}]\\ &\quad+\left(4c_{1}\alpha^{2}+\frac{L\alpha^{2}}{1-\beta}\right)\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right].\end{split}$ (19)
In the rest of the proof, let us show that the sum of the last two terms in equation 19 is non-positive.
First of all, by Lemma 2 we know that
$\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]\leq\mathbb{E}\left[p^{k}_{\max}\sum_{i=1}^{k-1}a_{k,i}\|x^{i+1}-x^{i}\|^{2}\right],$
where
$a_{k,i}=L^{2}\beta^{k-i}\left(k-i+\frac{\beta}{1-\beta}\right).$
Or equivalently,
$\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]\leq\mathbb{E}\left[\sum_{i=1}^{k-1}p^{k}_{\max}a_{k,k-i}\|x^{k+1-i}-x^{k-i}\|^{2}\right],$
where
$a_{k,k-i}=L^{2}\beta^{i}\left(i+\frac{\beta}{1-\beta}\right).$
Therefore, to make the sum of the last two terms of equation 19 to be non-positive, we need to have
$c_{i+1}$ $\leq c_{i}-\left(4c_{1}\alpha^{2}+\frac{L\alpha^{2}}{1-\beta}\right)L^{2}p^{i}_{\max}\beta^{i}\left(i+\frac{\beta}{1-\beta}\right)$
for all $i\geq 1$. To satisfy this inequality, we choose
$c_{i+1}$ $=c_{i}-\left(4c_{1}\alpha^{2}+\frac{L\alpha^{2}}{1-\beta}\right)L^{2}\beta^{i}p^{i}_{\max}\left(i+\frac{\beta}{1-\beta}\right)$
for all $i\geq 1$, which implies that
$c_{i}$ $=c_{1}-\left(4c_{1}\alpha^{2}+\frac{L\alpha^{2}}{1-\beta}\right)L^{2}\sum_{l=1}^{i-1}\beta^{i}p^{i}_{\max}\left(i+\frac{\beta}{1-\beta}\right).$
To have $c_{i}>0$ for all $i\geq 1$, we can set $c_{1}$ as
$c_{1}$ $=\left(4c_{1}\alpha^{2}+\frac{L\alpha^{2}}{1-\beta}\right)L^{2}\hat{p}^{k}_{\max}\sum_{i=1}^{\infty}\beta^{i}\left(i+\frac{\beta}{1-\beta}\right).$
where, $\hat{p}^{k}_{\max}=\max_{i\in[k]}\{p^{i}_{\max}\}$. Since
$\sum_{i=1}^{j}i\beta^{i}=\frac{1}{1-\beta}\left(\frac{\beta(1-\beta^{j})}{1-\beta}-j\beta^{j+1}\right),$
we have $\sum_{i=1}^{\infty}i\beta^{i}=\frac{\beta}{(1-\beta)^{2}}$ and
$c_{1}=\left(4c_{1}\alpha^{2}+\frac{L\alpha^{2}}{1-\beta}\right)L^{2}\hat{p}^{k}_{\max}\frac{\beta+\beta^{2}}{(1-\beta)^{2}},$
which implies that
$c_{1}=\frac{\alpha^{2}L^{3}\hat{p}^{k}_{\max}\frac{\beta+\beta^{2}}{(1-\beta)^{3}}}{1-4\alpha^{2}\frac{\beta+\beta^{2}}{(1-\beta)^{2}}\hat{p}^{k}_{\max}L^{2}}.$ (20)
Notice that $\alpha\leq\frac{1-\beta}{2\sqrt{2}L\sqrt{\hat{p}^{k}_{\max}}\sqrt{\beta+\beta^{2}}}$ ensures $c_{1}>0$.
Therefore,
$\mathbb{E}[L^{k+1}-L^{k}]$ $\leq\left(-\alpha+\frac{3-\beta+2\beta^{2}}{2(1-\beta)}L\alpha^{2}+5c_{1}\alpha^{2}\right)\mathbb{E}[\|g^{k}\|^{2}]$
$\quad+\left(\frac{\beta^{2}}{2(1+\beta)}L\alpha^{2}\sigma^{2}+\frac{1}{2}L\alpha^{2}\sigma^{2}+c_{1}\alpha^{2}\sigma^{2}\left(2\frac{1-\beta}{1+\beta}+1-p^{k}_{\min}\right)\right).$
By telescoping equation 16, we obtain the convergence bound of our proposed algorithm under nonconvex settings.
Theorem 2.
Let Assumption 1 hold and $\alpha^{k}=\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$. Then, the iterates of Algorithm 3 satisfy
$\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}[\|g^{i}\|^{2}]$ $\leq\mathcal{O}\left(\frac{f(x^{1})-f^{*}}{k\alpha}+L\alpha\sigma^{2}\left(1+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{(1-\beta)}\right)\right),$ (21)
where $\bar{p}^{k}_{\min}=\frac{1}{k}\sum_{i=1}^{k}\bar{p}^{i}_{\min}$ and $\hat{p}^{k}_{\max}=\max_{i\in[k]}\{p^{i}_{\max}\}$.
Proof.
From equation 16 we know that
$\begin{split}\mathbb{E}[L^{k+1}-L^{k}]\leq-R_{1}\mathbb{E}[\|g^{k}\|^{2}]+R^{k}_{2},\end{split}$ (22)
where
$R_{1}$ $=-\alpha+\frac{3-\beta+\beta^{2}}{2(1-\beta)}L\alpha^{2}+4c_{1}\alpha^{2},$
$R_{2}$ $=\frac{\beta^{2}}{2(1+\beta)}L\alpha^{2}\sigma^{2}+\frac{1}{2}L\alpha^{2}\sigma^{2}+c_{1}\alpha^{2}\sigma^{2}\left(2\frac{1-\beta}{1+\beta}+1-p^{k}_{\min}\right).$
We further define
$\bar{R}_{2}$ $=\frac{\beta^{2}}{2(1+\beta)}L\alpha^{2}\sigma^{2}+\frac{1}{2}L\alpha^{2}\sigma^{2}+c_{1}\alpha^{2}\sigma^{2}\left(2\frac{1-\beta}{1+\beta}+1-\bar{p}^{k}_{\min}\right),$
where $\bar{p}^{k}_{\min}=\frac{1}{k}\sum_{i=1}^{k}\bar{p}^{i}_{\min}.$
Telescoping equation 22 yields
$L^{1}\geq\mathbb{E}[L^{1}-L^{k+1}]\geq R_{1}\sum_{i=1}^{k}\mathbb{E}[\|g^{i}\|^{2}]-\sum_{k=1}^{k}R^{k}_{2},$
and therefore
$\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}[\|g^{i}\|^{2}]\leq\frac{L^{1}}{kR_{1}}+\frac{\bar{R}_{2}}{R_{1}}.$ (23)
In the rest of the proof, we will appropriately bound $R_{1}$ and $\bar{R}_{2}$.
First, let us show that $R_{1}\geq\frac{\alpha}{2}$ and $\alpha\leq\min\left\{\frac{1-\beta}{L(4-\beta+\beta^{2})},\frac{1-\beta}{2\sqrt{2}L\sqrt{\hat{p}^{k}_{\max}}\sqrt{\beta+\beta^{2}}}\right\}$.
From equation 20 we know that
$c_{1}=\frac{\alpha^{2}L^{3}\hat{p}^{k}_{\max}\frac{\beta+\beta^{2}}{(1-\beta)^{3}}}{1-4\alpha^{2}\frac{\beta+\beta^{2}}{(1-\beta)^{2}}L^{2}\hat{p}^{k}_{\max}}.$
Since $\alpha\leq\frac{1-\beta}{2\sqrt{2}L\sqrt{\hat{p}^{k}_{\max}}\sqrt{\beta+\beta^{2}}}$, we have
$4\alpha^{2}\frac{\beta+\beta^{2}}{(1-\beta)^{2}}L^{2}\hat{p}^{k}_{\max}\leq\frac{1}{2}.$
Thus,
$c_{1}$ $\leq\alpha^{2}L^{3}\hat{p}^{k}_{\max}\frac{\beta+\beta^{2}}{(1-\beta)^{3}}\leq\frac{L}{8(1-\beta)}.$
Therefore, in order to ensure $R_{1}\geq\frac{\alpha}{2}$, it suffices to have
$\frac{3-\beta+\beta^{2}}{2(1-\beta)}L\alpha+\frac{\alpha L}{2(1-\beta)}\leq\frac{1}{2}$
which is equivalent to our condition $\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$.
For $\bar{R}_{2}$, we can upperbound $c_{1}$ using our condition $\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$. Thus,
$c_{1}$ $\leq\alpha^{2}L^{3}\hat{p}^{k}_{\max}\frac{\beta+\beta^{2}}{(1-\beta)^{3}}\leq\frac{\hat{p}^{k}_{\max}\beta L}{2(1-\beta)}.$
Therefore,
$\bar{R}_{2}$ $=\frac{\beta^{2}}{2(1+\beta)}L\alpha^{2}\sigma^{2}+\frac{1}{2}L\alpha^{2}\sigma^{2}+c_{1}\alpha^{2}\sigma^{2}\left(2\frac{1-\beta}{1+\beta}+1-\bar{p}^{k}_{\min}\right)$
$\leq\frac{\beta^{2}}{2(1+\beta)}L\alpha^{2}\sigma^{2}+\frac{1}{2}L\alpha^{2}\sigma^{2}+\frac{\hat{p}^{k}_{\max}\beta L\alpha^{2}\sigma^{2}}{(1+\beta)}+L\alpha^{2}\sigma^{2}\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\frac{\beta}{1-\beta}$
$\leq\left(\frac{2\beta^{2}+8\hat{p}^{k}_{\max}}{2(1+\beta)}+\frac{1}{2}+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{8(1-\beta)}\right)L\alpha^{2}\sigma^{2}.$
By putting them all together, we obtain
$\begin{split}\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}[\|g^{i}\|^{2}]&\leq\frac{2\left(f(x^{1})-f^{*}\right)}{k\alpha}+\left(\frac{2\beta^{2}+8\hat{p}^{k}_{\max}}{2(1+\beta)}+\frac{1}{2}+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{8(1-\beta)}\right)L\alpha\sigma^{2}\\ &=\mathcal{O}\left(\frac{f(x^{1})-f^{*}}{k\alpha}+L\alpha\sigma^{2}\left(1+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{(1-\beta)}\right)\right).\end{split}$

### v2

@@
-Philip ${}^{*\,\sharp}$ Yandex, MIPT &Aleksandr Beznosikov
+FRUGAL: Memory-Efficient Optimization by Reducing State Overhead for Scalable Training
 Abstract
-With the increase in the number of parameters in large language models, the process of pre-training and fine-tuning increasingly demands larger volumes of GPU memory.
+With the increase in the number of parameters in large language models, the training process increasingly demands larger volumes of GPU memory.
 A significant portion of this memory is typically consumed by the optimizer state.
-To overcome this challenge, recent approaches such as low-rank adaptation (LoRA (Hu et al., 2021)), low-rank gradient projection (GaLore (Zhao et al., 2024a)), and blockwise optimization (BAdam (Luo et al., 2024)) have been proposed.
+To overcome this challenge, recent approaches such as low-rank adaptation (LoRA), low-rank gradient projection (GaLore), and blockwise optimization (BAdam) have been proposed.
 However, in all these algorithms, the effective rank of the weight updates remains low-rank, which can lead to a substantial loss of information from the gradient.
@@
 In this paper, we introduce FRUGAL (Full-Rank Updates with GrAdient spLitting), a new memory-efficient optimization framework.
-FRUGAL leverages gradient splitting to perform low-dimensional updates using advanced algorithms (such as Adam), while updates along the remaining directions are executed via state-free methods like SGD or signSGD (Bernstein et al., 2018).
+FRUGAL leverages gradient splitting to perform low-dimensional updates using advanced algorithms (such as Adam), while updates along the remaining directions are executed via state-free methods like SGD or signSGD.
 Our framework can be integrated with various low-rank update selection techniques, including GaLore and BAdam.
 We provide theoretical convergence guarantees for our framework when using SGDM for low-dimensional updates and SGD for state-free updates.
-Additionally, our method consistently outperforms concurrent approaches across various fixed memory budgets, achieving state-of-the-art results in pre-training and fine-tuning tasks while balancing memory efficiency and performance metrics.
-\$*\$\$*\$footnotetext: Work was mostly done during the internship at Mohamed bin Zayed University of Artificial Intelligence.$\sharp$\$$\sharp$\$footnotetext: The corresponding author: zmushko.ph.a@gmail.com\$$\diamond$\$$\diamond$\$footnotetext: Moscow Institute of Physics and Technology §$\S$ Mohamed bin Zayed University of Artificial Intelligence\$$\dagger$\$$\dagger$\$footnotetext: Ivannikov Institute for System Programming of the Russian Academy of Sciences $\ddagger$ Innopolis University
+Additionally, our method consistently outperforms concurrent approaches, achieving state-of-the-art results in pre-training and fine-tuning tasks while balancing memory efficiency and performance metrics.
+Machine Learning, ICML
 1 Introduction
-In recent years, Large Language Models (LLMs) such as GPT (OpenAI, 2023) and LLaMA-3 Dubey et al.
-(2024) have demonstrated remarkable performance across various disciplines (Brown, 2020; Yang et al., 2024; Romera-Paredes et al., 2024).
+In recent years, Large Language Models (LLMs) such as GPT (OpenAI, 2023) and LLaMA-3 (Dubey et al., 2024) have demonstrated remarkable performance across various disciplines (Brown, 2020; Yang et al., 2024; Romera-Paredes et al., 2024).
 However, a critical factor in achieving these results is the size of these models (Hoffmann et al., 2022).
-A larger number of parameters not only increases computational cost but also significantly raises memory requirements.
-For instance, training an 8 billion parameter LLaMA model in a 16-bit format necessitates each parameter to occupy 2 bytes, resulting in 16GB for storing the parameters and an additional 16GB for gradients.
-Utilizing the Adam optimizer (Kingma, 2014), which is standard for pre-training and fine-tuning LLMs, adds a further 32GB of memory to store the $m$ and $v$ statistics, resulting in 64GB total amount of memory.
-Furthermore, to achieve higher-quality results, training in pure 16-bit format is often insufficient (Zamirai et al., 2020).
-This necessitates storing master weights and optimizer statistics in 32-bit format, leading to total memory demands that exceed the capacity of cutting-edge graphics cards, such as the A100-80GB.
+Increasing the number of parameters leads to higher computational and memory costs.
+For example, an 8 billion parameter LLaMA-3 model in 16-bit format requires 32GB just for parameters and gradients.
+Using the standard Adam optimizer (Kingma, 2014) adds another 32GB for $m$ and $v$ statistics.
+Moreover, achieving high-quality results often requires 32-bit precision for weights and optimizer states (Zamirai et al., 2020), pushing memory requirements beyond even high-end GPUs like the A100-80GB.
 Numerous research projects have been aimed at reducing these significant costs.
-These approaches include engineering solutions like gradient checkpointing Chen et al.
-(2016) and memory offloading (Rajbhandari et al., 2020), which do not change the training trajectory.
-There are also methods that adjust the training algorithm by decreasing the number of trainable parameters (Frankle & Carbin, 2018; Wang et al., 2023; Sreenivasan et al., 2022; Horváth et al., 2024) or their bit precision (Wortsman et al., 2023), as well as optimizer statistics (Dettmers et al., 2021; Shazeer & Stern, 2018; Zhang et al., 2024c).
-In this work, we concentrate on the latter category.
-Parameter-Efficient Fine-Tuning (PEFT) methods, such as LoRA (Hu et al., 2021), Dora (Liu et al., 2024), and BitFit (Zaken et al., 2021) reduce memory costs by training a relatively small number of parameters compared to the size of the original model, while the remaining modules are frozen.
-This approach has proven effective for the task of efficient fine-tuning of pre-trained language models.
-However, PEFT methods have a fundamental limitation: parameter updates always lie in a low-dimensional subspace, which prevents the use of these methods for pre-training (Lialin et al., 2023) and may constrain their capabilities in fine-tuning (Zhang et al., 2024a).
-Recent works, such as GaLore (Zhao et al., 2024a), ReLoRA (Lialin et al., 2023), and BAdam (Luo et al., 2024), offer a solution to this problem.
-These methods enable higher-dimensional full-parameter learning by periodically changing the optimizable low-rank subspace, which we will call the state-full subspace.
+These approaches include engineering solutions like gradient checkpointing (Chen et al., 2016) and memory offloading (Rajbhandari et al., 2020), which do not change the training trajectory.
+There are also methods that adjust the training algorithm by decreasing the number of trainable parameters (Frankle & Carbin, 2018; Horváth et al., 2024) or their bit precision (Wortsman et al., 2023), as well as optimizer statistics (Dettmers et al., 2021; Shazeer & Stern, 2018; Zhang et al., 2024c).
+Parameter-Efficient Fine-Tuning (PEFT) methods, such as LoRA (Hu et al., 2021) and Dora (Liu et al., 2024b) reduce memory costs by training a relatively small number of parameters compared to the size of the original model, while the remaining modules are frozen.
+This approach has proven effective for the task of efficient fine-tuning of pre-trained models.
+However, PEFT methods have a fundamental limitation: parameter updates always lie in a low-dimensional subspace $L$, which prevents the use of these methods for pre-training (Lialin et al., 2023) and may restrict their capabilities in fine-tuning (Zhang et al., 2024a).
+Recent works, such as GaLore (Zhao et al., 2024a), ReLoRA (Lialin et al., 2023) and BAdam (Luo et al., 2024) offer a solution to this problem.
+These methods enable higher-dimensional full-parameter learning by periodically changing the optimizable low-rank subspace $L$.
 However, even though these methods result in overall parameter changes that are high-dimensional, the updates in each step remain low-dimensional.
-The dimensionality of the state-free subspace significantly exceeds that of the state-full subspace.
+The dimensionality of the frozen subspace $\dim M=\dim L^{\perp}$ significantly exceeds $\dim L$.
 The remaining information contained in the gradient is not utilized for parameter updates.
@@
 Figure 1: FRUGAL reduces memory usage by splitting gradient updates into low-dimensional updates with advanced optimizers (e.g., AdamW) and using state-free methods (e.g., SignSGD and SGD) for the rest.
+We present the FRUGAL framework, designed to bridge this gap.
+Our approach stems from a crucial observation: although memory constraints prevent using optimizers with auxiliary optimizer state — such as Adam (Kingma, 2014) — in the remaining subspace $M$, one still can update $M$ using state-free optimization algorithms like Stochastic Gradient Descent (SGD) or signSGD (Bernstein et al., 2018).
+This solution allows for high-dimensional updates, which provides additional opportunities to explore the parameter space and improves convergence.
+We will further refer to the subspaces $L$ and $M$ according to the types of optimizers used for their updates - state-full and state-free.
+Contributions. We summarize the main contributions of our work as follows:
+  • We present a new memory-efficient optimization framework that combines the use of advanced optimization algorithms for the state-full subspace with state-free algorithms for the complementary subspace.
+The framework supports various types of state-full optimizers, state-free optimizers, and different methods for projecting the gradient onto the state-full subspace.
+  • We provide theoretical convergence guarantees for our framework.
+In the proof, we consider the case with SGDM as the state-full optimizer and SGD as the state-free optimizer, and we show that FRUGAL matches the best-known convergence rate in many scenarios.
+  • To verify the practical applicability of FRUGAL, we conduct extensive experiments in popular real-world scenarios [^1] The code is available at https://anonymous.4open.science/r/FRUGAL-D3CA..
+In these experiments, we pre-train LLaMA-like models (up to 1B parameters) on the Colossal Clean Crawled Corpus (C4) dataset (Raffel et al., 2020) and fine-tune RoBERTa (Liu, 2019) on the GLUE benchmark (Wang, 2018).
+The results show that our method significantly outperforms previous memory-efficient algorithms while using less memory budget.
+  • We demonstrate that only the Output layer in transformer-like models requires advanced optimizers like Adam, while other modules (including RMSNorms and Embeddings) can use simpler methods like signSGD without significant performance loss.
+This opens up new possibilities for memory-efficient training and provides crucial insights into Transformers learning dynamics.
 Algorithm 1 FRUGAL(State-Full, State-Free)
-Input: model $f_{\mathbf{\theta}}(\cdot)$ with $p$ parameters $\{\theta_{i}\in\mathbb{R}^{d_{i}}\}_{i=1}^{p}$, loss function $\mathcal{L}$, gradient projectors $P_{k,i}$ for $i\in[p],k\in[K]$.
+Input: model $f_{\mathbf{\theta}}$ with $p$ parameter sets $\{\theta_{i}\in\mathbb{R}^{d_{i}}\}_{i=1}^{p}$, loss $\mathcal{L}$, gradient projectors $\{P_{k,i}\}_{i=1}^{p},$ number of steps $K$
 1:  for $k=1,2,\ldots K$ do
 2:     get data batch $(x,y)$
-3:     compute $\ell\leftarrow\mathcal{L}(f_{\theta}(x),y)$triangleright Forward
+3:     compute $\ell\leftarrow\mathcal{L}(f_{\theta}(x),y)$Forward}
 4:     for $g_{i}=\frac{\partial\ell}{\partial\theta_{i}}$ from Backward do
-5:         $g_{\text{full},i}\leftarrow P_{k,i}(g_{i}),\;$triangleright Project Grad
-6:         $g_{\text{free, i}}\leftarrow g_{i}-P^{-1}_{k,i}(g_{\text{full},i})$triangleright Residual
-7:         $\texttt{s}_{\theta_{i}}\leftarrow[P_{k,i}(P^{-1}_{k-1,i}(s),\;s\in\texttt{s}_{\theta_{i}}]$$\triangleright$ Project States
+5:        $g_{\text{full},i}\leftarrow P_{k,i}(g_{i}),\;$Project Grad}
+6:        $g_{\text{free, i}}\leftarrow g_{i}-P^{-1}_{k,i}(g_{\text{full},i})$Residual}
+7:        $\texttt{s}_{\theta_{i}}\leftarrow[P_{k,i}(P^{-1}_{k-1,i}(s),\;s\in\texttt{s}_{\theta_{i}}]$Project state}
 8:        $u_{\text{full, i}}\leftarrow\texttt{State-Full.update}(\theta_{i},g_{\text{full},i},\texttt{s}_{\theta_{i}})$
@@
 12:  end for
-We present the FRUGAL framework, designed to bridge this gap.
-Our approach stems from a crucial observation: although memory constraints prevent the use of advanced optimizers like Adam (Kingma, 2014) in the state-free subspace, one still can update this subspace using state-free optimization algorithms like Stochastic Gradient Descent (SGD) or signSGD (Bernstein et al., 2018).
-This solution allows for high-dimensional updates, which provides additional opportunities to explore the parameter space and improves convergence.
-Contributions. We summarize the main contributions of our work as follows:
-  • We present a new memory-efficient optimization framework that combines the use of advanced optimization algorithms for the state-full subspace with state-free algorithms for the complementary subspace.
-The framework supports various types of state-full optimizers, state-free optimizers, and different methods for projecting the gradient onto the state-full subspace.
-  • We provide theoretical convergence guarantees for our framework.
-In the proof, we consider the case where SGDM acts as the state-full optimizer and SGD as the state-free optimizer, and we show that FRUGAL matches the best-known convergence rate in many scenarios.
-  • To verify the practical applicability of FRUGAL, we conduct extensive experiments in popular real-world scenarios.
-[^1] The code is available at https://github.com/fzmushko/FRUGAL.
-In these experiments, we pre-train LLaMA-like models (up to 1B parameters) on the Colossal Clean Crawled Corpus (C4) dataset (Raffel et al., 2020) and fine-tune RoBERTa (Liu, 2019) on the GLUE benchmark (Wang, 2018).
-The results show that our method significantly outperforms previous memory-efficient algorithms while using the same memory budget.
-  • We demonstrate that only the Logits layer in transformer-like models requires advanced optimizers like Adam, while other modules (including Embeddings and RMSNorms) can use simpler methods like signSGD without significant performance loss.
-This opens new possibilities for memory-efficient training and provides crucial insights into the learning dynamics of Transformers.
 2 Related work
 Memory-efficient full-parameter learning.
-Recent research has focused on reducing the memory footprint of LLMs by decreasing the size of optimizer states while maintaining their performance.
-Low-rank adaptation methods, such as LoRA (Hu et al., 2021), inject trainable rank decomposition matrices into each layer of the model, reducing memory requirements by optimizing only a few learnable adapters.
-ReLora (Lialin et al., 2023) builds upon this by merging low-rank adaptations into the main model weights during training, potentially increasing the total rank of the update.
+Recent research has focused on reducing the memory footprint of LLM by decreasing the size of the optimizer states while maintaining their performance.
+Low-rank adaptation methods, such as LoRA (Hu et al., 2021), inject trainable rank decomposition matrices into linear layers, reducing memory requirements by optimizing only a few learnable adapters.
+ReLora (Lialin et al., 2023) builds upon this by merging low-rank adaptations into the main model weights during training, increasing the total rank of the update.
 BAdam (Luo et al., 2024) leverages Block Coordinate Descent for full-parameter training by switching active blocks during fine-tuning.
-MicroAdam (Modoranu et al., 2024) compresses gradient information before feeding it into the optimizer state, significantly reducing memory footprint while enabling full parameter learning through error feedback mechanisms.
-GaLore (Zhao et al., 2024a) maintains full parameter learning by projecting gradients onto a low-rank subspace using truncated SVD decomposition, storing optimizer states in this reduced space.
-Notably, GaLore achieves good results in pre-training, with performance close to that of Adam.
+MicroAdam (Modoranu et al., 2024) compresses gradient information before feeding it into the optimizer state, significantly reducing the memory footprint while enabling full parameter learning with error feedback mechanisms.
+GaLore (Zhao et al., 2024a) maintains full parameter learning by projecting gradients onto a low-rank subspace using SVD decomposition, storing optimizer states in this reduced space.
 However, while these methods effectively reduce memory overhead, they all perform low-rank updates at each iteration.
-In contrast, our approach utilizes all available gradient information to perform full-dimensional updates at each optimizer step, offering a novel perspective on memory-efficient optimization for LLMs.
+In contrast, our approach utilizes all available gradient information to perform full-dimensional updates at each optimizer step, offering a novel perspective on memory-efficient optimization for LLM.
+However, we note that there also exist several concurrent works — Fira (Chen et al., 2024a), LDAdam (Robert et al., 2024), and Adamem (Vyas et al., 2024) — that also adopt full-dimensional updates for similar goals.
+See Appendix B for a comparison and detailed discussion.
 Other memory-efficient optimization.
@@
 (2021) and Li et al.
-(2024) decrease memory footprint by quantizing optimizer states to the lower-precision representations.
+(2024) decrease memory footprint by quantizing optimizer states to lower-precision representations.
 Lv et al.
-(2023) proposed to reduce weight gradient memory by fusing the backward operation with the optimizer update.
+(2023) proposed to reduce memory by fusing the backward operation with the optimizer update.
 Notably, these approaches are orthogonal to our method FRUGAL and can be combined with it for further memory efficiency.
 Block Coordinate Descent.
-Block Coordinate Descent (BCD) is a well-established optimization method with a rich history in mathematical optimization (Ortega & Rheinboldt, 2000; Tseng, 2001; Richtárik & Takáč, 2014; 2015b; Richtárik & Takác, 2016; Takáč et al., 2013; Richtárik & Takáč, 2015a).
+Block Coordinate Descent (BCD) is a well-established optimization method with a rich history in mathematical optimization (Ortega & Rheinboldt, 2000; Tseng, 2001; Richtárik & Takáč, 2014, 2015a, 2015b).
 In recent years, a specific instance of BCD, known as layer-wise learning, has been applied to deep learning.
-Notable examples include (Luo et al., 2024; Pan et al., 2024), which leverage this approach for LLM fine-tuning.
-To the best of our knowledge, our work presents the first theoretical analysis of an extended BCD framework (Section 4) where the remaining coordinates are also updated using a different algorithm.
-This novel approach extends traditional BCD techniques, opening new avenues for full model optimization in deep learning.
+Notable examples include Luo et al.
+(2024); Pan et al.
+(2024), which leverage this approach for LLM fine-tuning.
+To the best of our knowledge, our work presents the first theoretical analysis of an extended BCD framework (Section 5) where the remaining layers are also updated with a different algorithm.
+This novel approach extends traditional BCD techniques, opening new avenues for full model optimization.
+Figure 2: Histograms of principal angle cosines.
+The first three are taken between ${\bm{P}}_{t}$ and ${\bm{P}}_{t^{\prime}}$ from different iterations $t$ and $t^{\prime}$.
+${\bm{P}}$ is obtained from the truncated SVD decomposition of the gradient ${\bm{G}}$ of the Key projection from the 5th layer.
+The last histogram is taken between two random semi-orthogonal projections ${\bm{R}}$ and ${\bm{R}}^{\prime}$ for comparison.
 Sign-based methods for training language models.
@@
 (2023) and Zhao et al.
-(2024b) showed that signed descent with momentum reduces the performance gap with Adam.
-In contrast, to the best of our knowledge, we are the first to train the majority of language model’s parameters using signSGD without momentum, achieving minimal loss in quality.
-This approach further demonstrates the effectiveness of sign-based methods for LLM training, paving the way for more efficient and scalable optimization strategies in deep learning.
-Table 1: Comparison of different projection and optimization strategies on pre-training LLaMA-130M on C4 with Adam as the state-full algorithm.
-State-free subspace is updated for full-rank methods but frozen for low-rank methods.
-Low-rank SVD outperforms Low-rank Random at the beginning of training but falls behind later due to insufficient exploration of the parameter space.
-Method Validation perplexity after iterations $\downarrow$
-4k 20k 40k 100k 200k
-Low-rank SVD 39.75 26.35 24.38 22.30 21.11
-Low-rank Random 42.31 25.99 23.55 21.33 20.01
-Full-rank Random 37.26 23.46 21.53 19.66 18.64
-Full-rank SVD 33.96 22.54 21.01 19.30 18.35
-Full-rank RandK 36.38 23.02 21.25 19.70 18.63
-Full-rank Block 37.20 23.34 21.42 19.59 18.60
-Adam 33.95 21.90 20.56 18.97 18.13
-3 FRUGAL: Full-Rank Updates with GrAdient spLitting
-In this section, we first introduce two main building blocks of our method and then introduce our general framework.
+(2024b) showed that the signed descent with momentum reduces the performance gap with Adam.
+In contrast, to the best of our knowledge, we are the first to train the majority of language model parameters using signSGD without momentum, achieving minimal loss in quality.
+This approach further demonstrates the effectiveness of sign-based methods for LLM training, paving the way for more efficient and scalable optimization strategies.
+3 Empirical Analysis and Motivation
 3.1 The importance of exploring the entire space during the training process
 In recent work, Zhao et al.
-(2024a) proposed GaLore, an optimization method based on projecting the gradient matrix ${\bm{G}}$ of each layer onto a low-dimensional subspace.
-To obtain the projection matrix ${\bm{P}}$, they use SVD decomposition of ${\bm{G}}_{t}$, which is recomputed with frequency $T$.
+(2024a) proposed GaLore, an optimization method based on projecting the gradient matrix ${\bm{G}}$ of each Linear layer222 Since Linear layers contain most parameters and require most memory, we primarly focus on them.
+onto a low-dimensional subspace.
+To obtain the projection matrix ${\bm{P}}$, they use the SVD decomposition of ${\bm{G}}_{t}$, which is recomputed with frequency $T$.
 The vectors or rows of ${\bm{G}}$ are projected onto the first $r$ left or right singular vectors, respectively.
 This approach has theoretical foundations: the first $r$ singular vectors correspond to the first $r$ singular values and, therefore, should better utilize information from the spectrum of ${\bm{G}}$.
-The authors pointed out that calculating the SVD decomposition results in extra computational overhead, which can be as much as a 10% increase as the hidden size of the model grows.
-To minimize this cost and examine the significance of using SVD decomposition, one may wonder about the possibility of employing a random semi-orthogonal projection matrix ${\bm{R}}$ instead of projecting onto the first $r$ singular columns with ${\bm{P}}$.
-Surprisingly, we found that although SVD decomposition delivers an initial boost, subsequent training with random projection yields significant improvements.
-As an illustration, we took the pre-training [^2] See Section 5.1 for a detailed description and discussion on the experimental setup.
-of a 130M model with LLaMA-like architecture on the C4 dataset (Raffel et al., 2020).
-The results are presented in Table 1, where training with ${\bm{P}}$ corresponds to the row labeled “Low-rank SVD”, and training with ${\bm{R}}$ corresponds to the row labeled “Low-rank Random”.
-The ranks of both projections ${\bm{P}}$ and ${\bm{R}}$ equal $192$.
+Table 1: Comparison of different projection and state-free subspace optimization strategies on pre-training LLaMA-130M on C4 with AdamW as the state-full algorithm.
+Projection Optimizes state- Validation perplexity $\downarrow$
+type free subspace 4k 40k 200k
+SVD No 39.75 24.38 21.11
+Random No 42.31 23.55 20.01
+Random Yes 37.26 21.53 18.64
+SVD Yes 33.96 21.01 18.35
+RandK Yes 36.38 21.25 18.63
+Blockwise Yes 37.20 21.42 18.60
+AdamW 33.95 20.56 18.13
+Given the computational burden of SVD decomposition, a natural question arises about the possibility of employing a random semi-orthogonal projection matrix ${\bm{R}}$ as an alternative to projecting onto the first $r$ singular columns with ${\bm{P}}$.
+Surprisingly, while the SVD decomposition provides better initial performance, the random projection proves superiority in long-term training, yielding significant improvements.
+As an illustration, we took the pre-training [^3] See Section 6.1 for a detailed description and discussion.
+of a 130M model with LLaMA-like architecture on the C4 dataset.
+The results are presented in the first part of Table 1, where we compare SVD and Random projections.
 To investigate this phenomenon, we pre-trained the LLaMA-60M model and collected gradients ${\bm{G}}_{t}$ from different iterations $t$ for examination.
-Following the setup from GaLore (Zhao et al., 2024a), we computed SVD decompositions and extracted projections ${\bm{P}}_{t}$ with a rank of 128.
-We evaluated the similarity of the projection matrices by calculating the principal angles between different projections ${\bm{P}}_{t}$ at different steps.
-Similarly to the observations in Q-Galore (Zhang et al., 2024d), we found that these projections show minimal change during the training period; see Figure 2.
-Here, we take the projection matrix corresponding to the $5$-th layer and plot histograms of the cosine of the principal angles between pairs ${\bm{P}}_{t}$ and ${\bm{P}}_{t^{\prime}}$ from different iterations.
+We evaluated the similarity of the projection matrices by calculating the principal angles between the projections ${\bm{P}}_{t}$ from different steps.
+Similarly to the observations in Q-Galore (Zhang et al., 2024d), we found that these projections show minimal change during training; see Figure 2 for details.
+Here, we take the projection matrix of k_proj from $5$-th layer and plot histograms of the cosine of the principal angles between pairs ${\bm{P}}_{t}$ and ${\bm{P}}_{t^{\prime}}$ from different iterations.
 For comparison, we also include the random projections on the right.
-As can be seen, the distributions of cosines differ significantly for the ${\bm{P}}_{t}$ and for the random projections.
+As can be seen, the distributions of cosines differ significantly for ${\bm{P}}_{t}$ and for ${\bm{R}}_{t}$.
 While ${\bm{R}}_{t}$ feature no angles with cosines higher than $0.9,$ the top $57$ cosines for ${\bm{P}}_{t}$ surpass $0.9$, even for gradients $1000$ steps apart.
-This leads to the conclusion that although SVD decomposition generally better captures the information contained in the ${\bm{G}}_{t}$, the original GaLore algorithm updates weights only in a small subspace.
+This leads to the conclusion that although the SVD decomposition generally better captures the information contained in ${\bm{G}}_{t}$, the original GaLore algorithm updates the weights only in a small subspace.
 We hypothesize that training with random projections yields superior results due to the more extensive investigation of the optimizable space during the training process.
-This finding indicates that to achieve better convergence, it is important to seek out optimization algorithms that explore the entire space during the training process.
+This finding indicates that to achieve better convergence, it is important to find optimization algorithms that explore the entire space during the training process.
 3.2 Advantage of the Full-Rank Updates
-The insight from the Section 3.1 suggests that the training of language models performs significantly better when the entire parameter space is utilized during the training process.
-Given the importance of updating parameters in all directions, this poses the question: Is it optimal to use low-rank updates, as employed by methods such as GaLore (Zhao et al., 2024a), ReLoRA (Lialin et al., 2023), and BAdam (Luo et al., 2024)?
-Using low-rank updates means the effective rank of the update is significantly smaller than the full dimensionality of the parameter space, inevitably leading to a loss of valuable information contained in the gradient.
+The insight from Section 3.1 suggests that the training of language models performs significantly better when the entire parameter space is explored during the training process.
+Given the importance of updating parameters in all directions, this poses the question: Is it optimal to use low-rank updates, as employed by methods such as GaLore, ReLoRA, and BAdam?
+The effective rank of low-rank updates is significantly smaller than the full dimensionality of the parameter space, inevitably leading to a loss of valuable information contained in the gradient.
 However, the method to leverage the full-rank gradient for updating parameters is not readily obvious.
-Using algorithms like Adam (Kingma, 2014) is not an option due to the memory overhead they introduce, which is precisely what we aim to avoid.
+Using algorithms like Adam (Kingma, 2014) is not an option due to the memory overhead they introduce, which is exactly what we aim to avoid.
 An alternative approach is to use state-free optimizers such as SGD or signSGD (Bernstein et al., 2018).
-Unfortunately, SGD have been shown to be ineffective for training transformer models, as shown in Zhang et al.
+Unfortunately, SGD has been shown to be ineffective for training transformer models, as shown in Zhang et al.
 (2020); Pan & Li (2023).
-Figure 2: Histograms of principal angle cosines.
-The first three are taken between ${\bm{P}}_{t}$ and ${\bm{P}}_{t^{\prime}}$ from different iterations $t$ and $t^{\prime}$.
-${\bm{P}}$ is obtained from the truncated SVD decomposition of the gradient ${\bm{G}}$ of the Key projection from the 5th layer.
-The last histogram is taken between two random semi-orthogonal projections ${\bm{R}}$ and ${\bm{R}}^{\prime}$ for comparison.
 Nevertheless, a recent study Zhao et al.
-(2024b) suggests a promising methodology: while SGDM doesn’t generally work well with transformers, using SGDM for the majority of parameters and Adam for a selected subset can lead to effective training.
+(2024b) suggests a promising methodology: while SGDM generally does not work well with transformers, using SGDM for the majority of parameters and Adam for a selected subset can lead to effective training.
 This raises the question: Could a hybrid approach using SGD or signSGD instead of SGDM be viable?
 If the key subset of parameters is handled by advanced algorithms, can the other parameters be trained effectively with state-free optimizers?
-To address this question, we conducted an experiment on LLaMA-130m, where we utilized the Adam (Kingma, 2014) for state-full parameters and signSGD (Bernstein et al., 2018) for state-free parameters.
-A detailed description of the experimental setup can be found at Section A.1.
-The result is highlighted as the entry “Full-rank Random” in Table 1.
-Full-rank updates significantly enhance performance, approaching the efficiency of the memory-intensive Adam optimizer, which serves as a upper bound in terms of performance.
-These findings underscore the potential of state-free algorithms for updating a substantial portion of the parameter space, paving the way for efficient, scalable optimization methods that deliver high performance without the significant memory costs traditionally associated with state-of-the-art optimizers.
-3.3 FRUGAL as Optimization Framework
-The setup outlined at the conclusion of the Section 3.2 results in a general framework for memory-efficient optimization.
-Briefly, it operates as follows: the entire space is partitioned into state-full and state-free subspaces.
-The state-full subspace is updated using advanced algorithms, while the state-free subspace is updated using a state-free method.
+To address this question, we conducted an experiment on LLaMA-130M, where we utilized the Adam (Kingma, 2014) for state-full parameters and signSGD (Bernstein et al., 2018) for state-free parameters [^4] See detailed description of the setup in Section A.1.
+Once again we used Random projection and highlighted the result in the second part of Table 1.
+Full-rank updates significantly enhance performance, approaching the efficiency of the memory-intensive Adam optimizer.
+These findings underscore the potential of state-free algorithms for updating a substantial portion of the parameter space, paving the way for efficient and scalable optimization methods that deliver high performance without the significant memory costs traditionally associated with state-of-the-art optimizers.
+4 Full-Rank Updates with GrAdient spLitting
+General framework.
+The setup outlined in the conclusion of Section 3.2 results in a general framework for memory-efficient optimization.
+It operates as follows: the entire space is partitioned into state-full and state-free subspaces.
+The state-full subspace is updated using an advanced algorithm, while the state-free subspace is updated using a state-free method.
 After a certain number of steps, the state-full subspace is changed to better explore the optimization space.
+A formal description is presented in Algorithm 1.
+We note that this framework allows for variation not only in the state-full optimizer but also in the choice of projection and state-free optimizer.
 However, determining the optimal state-free optimizer and the projection method onto the state-full subspace is not readily apparent.
@@
 State-free optimizer.
-We conducted a preliminary experiment updating all parameters using state-free algorithms to choose between SGD and signSGD (Bernstein et al., 2018).
-Table 8 presents these results.
-After testing various learning rates, we found that signSGD consistently outperforms SGD, leading us to favor signSGD.
+We conducted a preliminary experiment using different state-free algorithms to choose between SGD and signSGD (Bernstein et al., 2018).
+Table 10 shows that signSGD outperforms SGD, leading us to favor signSGD.
 We attribute this performance to the similarities between signSGD and Adam (Kingma, 2014), as noted in Balles & Hennig (2018); Balles et al.
@@
 When selecting a projection method, it is crucial to strike a balance between quality and memory efficiency.
-For instance, the approach using the projection matrix from SVD decomposition, as in GaLore (Zhao et al., 2024a), better preserves the information embedded in the gradient but requires additional memory for storing projection matrices and computational resources for performing the SVD.
-To reduce computational demands, one could employ random coordinate projection denoted as RandK, but this requires additional memory or recomputation [^3] See Appendix B for discussion on the memory requirements for different projection methods..
-A more structured alternative is to select not random matrix elements but entire random columns or rows.
+When using SVD decomposition for projection matrices, as in GaLore (Zhao et al., 2024a), the method better preserves the information embedded in the gradient but requires additional memory for storing projection matrices and computational resources for performing the SVD.
+To reduce computational demands, one could employ random coordinate projection denoted as RandK, but this requires additional memory or recomputation [^5] See Appendix C for discussion on the memory requirements for different projection methods..
+A more structured alternative is to select not random entries but entire random columns or rows.
 The most aggressive approach follows the method from BAdam, wherein an entire block is chosen as the state-full subspace.
-The performance results obtained with all these variants are presented in Table 1.
-Full-rank SVD outperforms both Full-rank RandK and Full-rank Block methods, demonstrating comparable performance.
-The superior performance of SVD projections can be explained by their ability to extract the principal information from the gradient.
+The performance results obtained with all these variants are presented in the second part of Table 1.
+SVD slightly outperforms both RandK and Block projections, demonstrating comparable performance.
 Nonetheless, a downside is the increased compute and memory demand from SVD.
 Therefore, we opt for the blockwise selection, as it is the most memory-efficient — requiring only the storage of active block indices.
-A formal description of the final algorithm is presented in Algorithm 1.
-It’s worth noting that this framework allows for variation not only in the State-Full algorithm but also in the choice of projection and State-Free optimizer.
-In our experiments in Section 5, we use a specific variant with Adam as the State-Full optimizer, signSGD as the State-Free optimizer, and blockwise projection.
+In experiments in Section 6, we use a specific variant with AdamW as the State-Full optimizer and signSGD as the State-Free optimizer.
+We primarily employ blockwise projection but switch to column-wise projection when the number of parameters in any single block exceeds memory budget, as detailed in Section 7.
+In addition, PyTorch-like pseudocode of our framework is presented in Appendix G.
 For Line 7, state projection, in Algorithm 1, we note that if the projection does not change, i.e., $P_{k,i}=P_{k-1,i}$, then $P_{k,i}(P^{-1}_{k-1,i}(s))=s$.
@@
 However, either resetting or projecting states is important since we want projected gradients and optimizer states to reside in the same space.
-For instance, GaLore ignores this step, which leads to degraded performance when projections are updated frequently; see Appendix C for details.
+For instance, GaLore ignores this step, which leads to degraded performance when projections are updated frequently; see Appendix D and Section 6.4 for details.
 Algorithm 2 FRUGAL(SGDM, SGD)
@@
 1:  for $k=1,2,\ldots$ do
-2:     Compute stochastic gradient $\tilde{g}^{k}\leftarrow\nabla f_{\zeta^{k}}(x^{k})$;
-3:     Update momentum vector $\tilde{m}_{j}^{k}\leftarrow(1-\beta)\tilde{g}_{j}^{k}+\beta\begin{cases}\tilde{m}_{j}^{k-1}&\text{if }j\in J_{k},\\ 0&\text{otherwise;}\end{cases}$
-4:     Compute update vector $\tilde{u}_{j}^{k}\leftarrow\begin{cases}\tilde{m}_{j}^{k}&\text{if }j\in J_{k},\\ \tilde{g}_{j}^{k}&\text{otherwise;}\end{cases}$
-5:     Update iterate $x^{k+1}\leftarrow x^{k}-\alpha^{k}\tilde{u}^{k}$;
+2:     $\tilde{g}^{k}\leftarrow\nabla f_{\zeta^{k}}(x^{k})$
+3:     $\tilde{m}_{j}^{k}\leftarrow(1-\beta)\tilde{g}_{j}^{k}+\beta\begin{cases}\tilde{m}_{j}^{k-1}&\text{if }j\in J_{k},\\ 0&\text{otherwise;}\end{cases}$
+4:     $\tilde{u}_{j}^{k}\leftarrow\begin{cases}\tilde{m}_{j}^{k}&\text{if }j\in J_{k},\\ \tilde{g}_{j}^{k}&\text{otherwise;}\end{cases}$
+5:     $x^{k+1}\leftarrow x^{k}-\alpha^{k}\tilde{u}^{k}$
 6:  end for
@@
 This special case of FRUGAL is provided in Algorithm 2.
-We minimize the following objective
-$\textstyle\min_{x\in\mathbb{R}^{d}}\big{\{}f(x):=\mathbb{E}_{\zeta}[f_{\zeta}(x)]\big{\}},$ (1)
-where we access $f$ via a stochastic oracle that takes $x$ as input and returns $(f_{\zeta}(x),\nabla f_{\zeta}(x))$.
+We minimize the objective
+$\textstyle\min_{x\in\mathbb{R}^{d}}\big{\{}f(x):=\mathbb{E}_{\zeta{{}^{k}}}[f_{\zeta{{}^{k}}}(x)]\big{\}},$ (1)
+where we access $f$ via a stochastic oracle that takes $x$ as input and returns $(f_{\zeta{{}^{k}}}(x),\nabla f_{\zeta{{}^{k}}}(x))$.
 5.1 Notation and Preliminaries
 […]
-Assumption 1.
+$\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|.$ (2)
+Table 2: Comparison of validation perplexity and memory estimation for various optimization methods across LLaMA model scales trained on C4.
+We also indicate the additional memory overhead introduced by the optimization algorithm.
+The values are calculated assuming that each float value occupies 4 bytes (float32).
+$\rho$ denotes the proportion of the Linear layer parameters in the state-full subspace.
+Note that Embeddings, RMSNorms, and Output layer are always trained with AdamW.
+60M 130M 350M 1B
+AdamW 22.73 (0.43G) 18.13 (1.00G) 14.43 (2.74G) 12.02 (9.98G)
+GaLore, $\rho=0.25$ 25.68 (0.30G) 21.11 (0.54G) 16.88 (1.10G) 13.69 (3.41G)
+BAdam, $\rho=0.25$ 24.86 (0.29G) 20.34 (0.52G) 16.41 (1.05G) 13.75 (3.23G)
+FRUGAL, $\rho=0.25$ 23.59 (0.29G) 18.60 (0.52G) 14.79 (1.05G) 12.32 (3.23G)
+FRUGAL, $\rho=0.0$ 24.06 (0.24G) 18.90 (0.37G) 15.03 (0.49G) 12.63 (0.98G)
+Training tokens 20B 20B 24B 30B
+Number of iterations 200k 200k 240k 300k
+Assumption 5.1.
 We make the following assumptions, which are standard in non-convex stochastic optimization; see (Liu et al., 2020).
   1. 1.
-    Smoothness: The objective $f(x)$ in equation 1 is $L$-smooth.
+    Smoothness: The objective $f(x)$ in equation 1 is $L$-smooth (Equation 2).
   2. 2.
@@
   3. 3.
-    Independent samples: The random samples $\{\zeta_{k}\}_{k=1}^{\infty}$ are independent.
+    Independent samples: The random samples $\{\zeta^{k}\}_{k=1}^{\infty}$ are independent.
   4. 4.
@@
 Below, we present the main convergence theorem.
-Theorem 1.
-Let Assumption 1 hold and $\alpha^{k}=\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$. Then, the iterates of Algorithm 2 satisfy
-$\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}[\|g^{i}\|^{2}]$ $=\mathcal{O}\left(\frac{f(x^{1})-f^{*}}{k\alpha}+L\alpha\sigma^{2}\left(1+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{(1-\beta)}\right)\right),$ (2)
+Theorem 5.2.
+Let Assumption 5.1 hold and $\alpha^{k}=\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$. Then, the iterates of Algorithm 2 satisfy
+$\textstyle\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}[\|g^{i}\|^{2}]$ $=\mathcal{O}\biggl{(}\frac{f(x^{1})-f^{*}}{k\alpha}+$
+$+L\alpha\sigma^{2}\Bigl{(}1+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{(1-\beta)}\Bigr{)}\biggr{)},$
 where $\bar{p}^{k}_{\min}=\frac{1}{k}\sum_{i=1}^{k}\bar{p}^{i}_{\min}$ and $\hat{p}^{k}_{\max}=\max_{i\in[k]}\{p^{i}_{\max}\}$.
@@
 In this case, we have $\bar{p}^{k}_{\min}=1$ for SGDM and $\hat{p}^{k}_{\max}=0$ for SGD.
-Therefore, the resulting rate is $\mathcal{O}\left(\nicefrac{{1}}{{k\alpha}}+L\alpha\sigma^{2}\right)$, which recovers the best-known rate for both SGD and SGDM under these assumptions Liu et al.
-(2020).
+Therefore, the resulting rate is $\mathcal{O}\left(\nicefrac{{1}}{{k\alpha}}+L\alpha\sigma^{2}\right)$, which recovers the best-known rate for both SGD and SGDM under these assumptions (Liu et al., 2020).
 Furthermore, if at each step each coordinate is sampled independently with probability $p$, we have $\bar{p}^{k}_{\min}=\hat{p}^{k}_{\max}=p$.
@@
 Thus, the rate becomes $\mathcal{O}\left(\nicefrac{{1}}{{k\alpha}}+\nicefrac{{L\alpha\sigma^{2}}}{{1-\beta}}\right)$, which is worse by a factor of $\nicefrac{{1}}{{1-\beta}}$.
-However, this is expected since the bias from momentum is not outweighed by the variance reduction effect, as only the coordinates with momentum enjoy reduced variance; see Lemmas 1 and 2 in the appendix for details.
-5 Experiments
-This section presents the main experimental results of the paper. To evaluate the performance of FRUGAL, we conducted experiments both on the pre-training and fine-tuning of language models.
-5.1 Pre-training experiments
+However, this is expected since the bias from momentum is not outweighed by the variance reduction effect, as only the coordinates with momentum enjoy reduced variance; see Lemmas E.2 and E.3 in the appendix for details.
+6 Pre-training experiments
+In this section, we evaluate the performance of FRUGAL on the language models pre-training.
+6.1 Comparison to existig memory-efficient algorithms
+To begin, we compare our framework with existing memory-efficient methods across four sizes of LLaMA-based architectures: 60M, 130M, 350M, and 1B.
 Setup.
-The core setup for pre-training is taken from the Zhao et al.
+The core setup for pre-training is taken from Zhao et al.
 (2024a).
-We utilize LLaMA-based (Touvron et al., 2023a) model architectures with up to 1B parameters and train them on the Colossal Clean Crawled Corpus (C4) dataset (Raffel et al., 2020).
+We utilize LLaMA-based (Touvron et al., 2023a) model architectures and train them on the Colossal Clean Crawled Corpus (C4) dataset (Raffel et al., 2020).
 The C4 dataset is intended for pre-training, making this setup a good approximation of real-world applications.
 A detailed description of the setup can be found in Section A.1.
-However, we made several modifications that we would like to discuss in detail below.
+Table 3: Perplexity and memory consumption (weights, gradients and optimizer states) of different size LLaMA models pre-trained on C4 for 100k iterations (10B tokens) using AdamW with pure bf16 of mixed precision.
+Model size Format Memory Perplexity
+175M Mixed Precision 2.0GB 17.43
+350M Pure bf16 2.1GB 17.75
+350M Mixed Precision 4.2GB 15.16
+1.3B Pure bf16 7.7GB 16.51
+However, we made several critical modifications compared to Zhao et al.
+(2024a) to align the experimental setup with practical training scenarios.
+Below, we discuss each modification and provide a detailed rationale for these decisions.
+Table 4: Perplexity of LLaMA-130M models pre-trained on C4 for 100k iterations (10B tokens).
+The leftmost column indicates the modules moved to the state-free set and trained using signSGD.
+The results show that Output layer, unlike Embeddings and RMSNorms, are exceptionally responsive to the choice of optimization algorithm from AdamW to signSGD.
+State-free modules Perplexity $\downarrow$
+[…]
+Linear, RMSNorms 20.07
+Linear, Embeddings 20.48
+Linear, Embeddings, RMSNorms 20.55
+Linear, Output layer 34.66
   • Training Duration.
 The training approach in Zhao et al.
-(2024a) aligns with the empirical rule from scaling laws (Hoffmann et al., 2022), which suggests using approximately 20 times the model size in tokens for training.
+(2024a) aligns with the empirical rule from scaling laws (Hoffmann et al., 2022), which suggests using approximately 20 times the size of the model in tokens for training.
 However, this number of tokens is far from achieving convergence.
 In practice, models are typically trained for significantly longer periods (Touvron et al., 2023b; Zhang et al., 2024b).
-One reason for this discrepancy is that the original scaling laws do not account for the inference of the model after training.
-Adjustments to scaling laws considering this parameter are discussed, for example, in (Sardana & Frankle, 2023).
-For our experiments we chose 200k steps for the 60M and 130M models, 240k for 350M model and 300k for the 1B model.
-  • Learning Rate.
-The authors of GaLore suggested using different learning rates for fixed unprojectable parameters (Embeddings, RMSNorms (Zhang & Sennrich, 2019), Logits) and the remaining projectable parameters (attention and FFN weights modules weights).
-However, introducing additional hyperparameters complicates the use of the algorithm.
-Since both sets of parameters are state-full and trained using the same optimization algorithm, we always used the same learning rate for them in FRUGAL and BAdam.
-For GaLore learning rate see Section 5.1.
-  • Mixed Precision instead of the pure bfloat16 training.
+One reason for this discrepancy is that the original scaling laws do not account for the inference of the model after training (Sardana & Frankle, 2023).
+For our experiments, we chose 200k steps for the 60M and 130M models, 240k for the 350M model, and 300k for the 1B and 3B models.
+  • Mixed Precision.
 Pure 16-bit training has been shown to potentially compromise model convergence and accuracy (Zamirai et al., 2020).
-This degradation stems from storing both the model weights and optimizer statistics in reduced precision formats such as float16 or bfloat16.
-However, these formats often lack sufficient precision in representing floating-point numbers.
-Consequently, mixed precision training has become a more common approach for training language models (Le Scao et al., 2023; Almazrouei et al., 2023)).
+This degradation occurs because formats such as float16 or bfloat16, used to store master weights, lack the numerical precision needed for accurate and fine-grained weight updates.
+Consequently, mixed precision training has become a more common approach for training language models (Le Scao et al., 2023; Almazrouei et al., 2023).
+Moreover, even when training with fp8, the master weights are typically stored in fp32 format (Liu et al., 2024a).
+    Our experimental results strongly support the importance of precision choice: in Table 3 we show that adopting pure bf16 training led to such significant performance degradation that doubling the model size failed to compensate for it, effectively negating any memory benefits from reduced precision storage.
 While training in pure 16-bit format is also possible, stochastic rounding (Gupta et al., 2015; Zamirai et al., 2020) is often employed to mitigate the aforementioned issue.
@@
 Hence, we primarily used Mixed Precision training for its illustrative value in understanding each method’s potential.
-However, for completeness, we also conducted experiments in pure bfloat16 format, detailed in our ablation study Section 5.1.2.
-Table 2: Comparison of validation perplexity and memory estimation for various optimization methods across LLaMA model scales trained on C4.
-Lower $\downarrow$ is better.
-In parentheses, we also indicate the additional memory overhead introduced by the optimization algorithm.
-The values are calculated assuming that each float value occupies 4 bytes (float32).
-Note that Embeddings, RMSNorms, and Logits are always trained using state-full Adam.
-60M 130M 350M 1B
-Adam 22.73 (0.43G) 18.13 (1.00G) 14.43 (2.74G) 12.02 (9.98G)
-GaLore, $\rho=0.25$ 25.68 (0.30G) 21.11 (0.54G) 16.88 (1.10G) 13.69 (3.41G)
-BAdam, $\rho=0.25$ 24.86 (0.29G) 20.34 (0.52G) 16.41 (1.05G) 13.75 (3.23G)
-FRUGAL, $\rho=0.25$ 23.59 (0.29G) 18.60 (0.52G) 14.79 (1.05G) 12.32 (3.23G)
-FRUGAL, $\rho=0.0$ 24.06 (0.24G) 18.90 (0.37G) 15.03 (0.49G) 12.63 (0.98G)
-Training tokens 20B 20B 24B 30B
-Number of iterations 200k 200k 240k 300k
-5.1.1 Comparison to existing memory-efficient algorithms
-To begin, we present the results of comparing FRUGAL with existing memory-efficient methods across four sizes of LLaMA-based architectures: 60M, 130M, 350M, and 1B [^4] See preliminary experimental results with LLaMA 7B in Appendix D.
-Baselines. We use the following methods as baselines for our approach:
-  • Full-rank Training. Training using memory-inefficient Adam. Weights, gradients, and statistics are stored and computed for all parameters. This serves as an upper bound for model performance.
+However, for completeness, we also conducted experiments in pure bfloat16 format, detailed in our ablation study Section 6.4.
+Baselines. We use the following methods as baselines:
+Table 5: Pre-training LLaMA 3B on C4 dataset for 300K steps. Validation perplexity for different iterations is reported.
+Method 100k 200k 300k
+AdamW 14.2 12.25 10.93
+FRUGAL, $\rho=0.25$ 14.33 12.42 11.07
+FRUGAL, $\rho=0.0$ 14.78 12.76 11.35
+Table 6: Evaluating FRUGAL for memory-efficient fine-tuning RoBERTa-Base on GLUE benchmark. Results represent the mean and standard deviation across 3 independent runs. Upper $\uparrow$ is better.
+Method Modules Rank CoLA STS-B MRPC RTE SST2 MNLI QNLI QQP Avg
+Full-parameter 63.6 91.2 90.2 78.7 94.8 87.6 92.8 91.9 86.4
+LoRA QV 8 63.8$\pm$.6 90.9$\pm$.1 89.1$\pm$.4 79.2$\pm$1.1 94.8$\pm$.2 87.6$\pm$.2 93.1$\pm$.1 90.6$\pm$.0 86.1
+GaLore All 8 60.0$\pm$.2 90.8$\pm$.1 89.0$\pm$.7 79.7$\pm$.9 94.9$\pm$.5 87.6$\pm$.1 93.3$\pm$.1 91.1$\pm$.1 85.8
+GaLore QV 8 56.1$\pm$.8 90.8$\pm$.2 88.1$\pm$.3 74.7$\pm$1.9 94.3$\pm$.1 86.6$\pm$.1 92.6$\pm$.1 89.4$\pm$.1 84.1
+FRUGAL QV 8 64.5$\pm$.7 91.1$\pm$.1 89.2$\pm$.3 82.4$\pm$.9 94.8$\pm$.2 87.4$\pm$.1 92.8$\pm$.1 91.4$\pm$.1 86.7
+FRUGAL None 0 64.8$\pm$.5 91.1$\pm$.1 89.1$\pm$.3 81.6$\pm$.6 94.9$\pm$.2 87.3$\pm$.1 92.8$\pm$.1 91.3$\pm$.1 86.6
+  • Full-rank Training.
+Training using memory-inefficient AdamW (Loshchilov, 2017).
+Weights, gradients, and statistics are stored and computed for all parameters.
+This serves as an upper bound for model performance.
   • GaLore.
@@
 (2024) proposed a block coordinate descent (BCD)-type optimization method termed BAdam.
-The parameters are divided into blocks, which are then updated one by one using Adam.
-Similar to GaLore, the optimized block is updated every $T$ steps.
+The parameters are divided into blocks, which are then updated one by one using AdamW.
+The optimized block is changed every $T$ steps.
 Although this method was initially proposed only for fine-tuning, it is the closest method to our FRUGAL.
@@
   • Other Algorithms.
-Among other relevant methods, ReLoRA (Lialin et al., 2023) and MicroAdam (Modoranu et al., 2024) can also be highlighted.
-However, we did not include them for comparison in this paper for the following reasons: 1.
-ReLoRA was evaluated in (Zhao et al., 2024a), where it significantly underperformed compared to GaLore with the same memory budget.
+Among other relevant methods, ReLoRA (Lialin et al., 2023), MicroAdam (Modoranu et al., 2024), Fira (Chen et al., 2024a), LDAdam (Robert et al., 2024), and Adamem (Vyas et al., 2024) can also be highlighted.
+However, we did not include them for comparison here for the following reasons: 1.
+ReLoRA: This method was evaluated in (Zhao et al., 2024a), where it significantly underperformed compared to GaLore.
 2.
-MicroAdam.
-Its current implementation only supports bfloat16 master weights, whereas our main experiments conducted with mixed precision.
+MicroAdam: Its current implementation only supports bfloat16 master weights, whereas our main experiments conducted with mixed precision.
+3.
+Fira, LDAdam, and Adamem: These methods are concurrent works that were published during the final stages of our work.
+Accordingly, the majority of our experiments were completed before we became aware of these recent developments.
 Main results.
 The results of our experiments are presented in Table 2, which includes both validation perplexity and memory footprint estimations for each method.
-We compared all memory-efficient methods under the same memory budget, specifically with a density $\rho=0.25$.
-Here, $\rho$ refers to the proportion of state-full parameters (see Section A.1 for a detailed explanation).
-We conducted a grid search to determine the optimal learning rate for Adam, which we then applied uniformly to FRUGAL and BAdam (Luo et al., 2024).
-For GaLore (Zhao et al., 2024a), we found that using this same learning rate produced better results than the rate originally suggested in their paper.
+We compared all memory-efficient methods under the same memory budget with a density $\rho=0.25$.
+Here, $\rho$ refers to the proportion of Linear layer parameters belonging to the state-full subspace.
+Similarly to GaLore, non-Linear modules (Embeddings, RMSNorms, Output layer) are optimized with AdamW.
+See Section A.1 for details.
+We conducted a grid search to determine the optimal learning rate for AdamW, which we then applied to FRUGAL and BAdam (Luo et al., 2024).
+For GaLore (Zhao et al., 2024a), we found that using this same learning rate produced better results than the originally suggested rate.
 This discrepancy might be attributed to our experiments involving a significantly larger number of training steps than those for which GaLore’s original learning rate was optimized.
-Table 2 demonstrates that FRUGAL significantly outperforms the memory-efficient baselines across all model sizes with the same memory budget, coming close to the performance of Adam.
-Table 3: Perplexity of LLaMA-130M models pre-trained on C4 for 100k iterations (10B tokens).
-The leftmost column indicates the modules moved to the state-free set and trained using signSGD.
-The results show that Logits, unlike Embeddings and RMSNorms, are exceptionally responsive to the choice of optimization algorithm from Adam to signSGD.
-State-free modules Perplexity $\downarrow$
-Linear (corresponds to the FRUGAL with $\rho=0.0$) 20.02
-Linear, RMSNorms 20.07
-Linear, Embeddings 20.48
-Linear, Embeddings, RMSNorms 20.55
-Linear, Logits 34.66
-Zero-density training.
-Table 2 also reveals a surprising result: FRUGAL with $\rho=0.0$ significantly outperforms both GaLore and BAdam, even when these competing methods use a higher density of $\rho=0.25$.
-Essentially, in this case, the parameters are divided into two parts — a state-full part consisting of the Embeddings, RMSNorms, and Logits, and a state-free part consisting of all other parameters.
+Table 2 demonstrates that FRUGAL significantly outperforms memory-efficient baselines across all model sizes with the same memory budget, coming close to the performance of AdamW.
+6.2 Zero-density training
+Table 2 also reveals a surprising result: FRUGAL with $\rho=0.0$ outperforms both GaLore and BAdam, even when these competing methods use a higher density of $\rho=0.25$.
+Essentially, for FRUGAL with $\rho=0.0$, the parameters are divided into two parts — a state-full part consisting of the Embeddings, RMSNorms, and Output layer, and a state-free part consisting of all other parameters.
 This division remains fixed throughout the training.
 We conducted additional experiments to determine the maximum subset of parameters that can be trained with a state-free optimizer without significant quality degradation.
-We systematically moved different combinations of the Embeddings, RMSNorms, and Logits from the state-full to the state-free set and observed the results during the training of LLaMA-130M.
-Table 3 reveals that the Logits demonstrates a dramatically higher sensitivity, with changes to its optimizer resulting in severe performance degradation.
-This finding aligns with results from (Zhao et al., 2024b), where the authors demonstrated that most parameters can be trained using SGDM, but the Logits require training with Adam.
-5.1.2 Ablation study
+We systematically moved different combinations of the Embeddings, RMSNorms, and Output layer from the state-full to the state-free set and observed the results during the training of LLaMA-130M.
+Table 4 reveals that the Output layer demonstrates a dramatically higher sensitivity, with changes to its optimizer resulting in severe performance degradation.
+This finding aligns with results from Zhao et al.
+(2024b), where the authors demonstrated that most parameters can be trained using SGDM, but the Output layer require training with AdamW.
+6.3 LLaMA 3B training
+To demonstrate the practical viability of our method for large-scale applications, we evaluated FRUGAL against AdamW on the pre-training of the LLaMA 3B model.
+Due to computational constraints, we conducted a single training run of 300k steps using a cosine learning rate scheduler with 10% warmup steps.
+We used a learning rate of 5e-4, weight decay of 0.1, and gradient clipping of 1.0, with other hyperparameters consistent with Section A.1.
+The results in Table 5 confirm that FRUGAL successfully scales to billion-parameter models without performance degradation, making it a viable option for industrial-scale applications.
+6.4 Ablation study
+Table 7: Accuracy $\uparrow$ comparison of memory-efficient methods on LLaMA 3.1-8B across 8 commonsense reasoning tasks.
+Method Rank BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg
+LoRA 32 73.39 88.19 79.48 95.13 83.50 90.24 79.52 84.00 84.18
+GaLore 32 74.13 88.36 78.97 94.88 83.82 91.08 80.2 85.80 84.65
+FRUGAL 32 74.25 88.47 79.12 95.15 86.11 91.37 79.86 84.80 84.89
+FRUGAL 0 71.44 88.47 80.40 94.74 86.35 90.95 81.14 86.00 84.94
 We also conducted additional experiments to verify the robustness of our framework to various hyperparameters.
-Firstly, an ablation study on the state-full subspace update frequency $T$ in Table 10 shows that the performance keeps improving up to $T=200$.
+First, we began by evaluating different model architectures.
+Experiments with GPT-2 124M (Radford et al., 2019) in Table 12 show that FRUGAL maintains its strong advantage over memory-efficient baselines, albeit with a somewhat wider gap to AdamW.
+Second, an ablation study on the state-full subspace update frequency $T$ in Table 14 shows that the performance keeps improving up to $T=200$.
 We note that, unlike in Zhao et al.
-(2024a), the perplexity does not significantly decrease even when reducing the update frequency to $T=10$ ($\sim 0.2$ drop vs.
+(2024a), the perplexity does not decrease significantly even when reducing the update frequency to $T=10$ ($\sim 0.2$ drop vs.
 $\sim 4.$ drop for GaLore).
-A detailed explanation for this result can be found in Appendix C.
-Second, when using other schedulers, the performance gap between FRUGAL and baselines remains consistent, as shown in Tables 6 and 6.
-Then, the results of training in pure bfloat16 are presented in Table 8, demonstrating consistency with our main experiments in Table 2, i.e., FRUGAL significantly outperforms the baselines across these variations.
+A detailed explanation for this result can be found in Appendix D.
+After that, when using other schedulers, the performance gap between FRUGAL and baselines remains consistent, as shown in Tables 16 and 16.
+Table 8 shows that the same holds for $\beta_{2}=0.95$ — another popular value for the second moment decay parameter in AdamW-like methods.
+FRUGAL also provides improvement over baselines for other state-full optimizers, as can be seen in experiments with Lion (Chen et al., 2024b) presented in Table 12.
+Then, the results of the training in pure bfloat16 are presented in Table 10, demonstrating consistency with our main experiments in Table 2, i.e., FRUGAL significantly outperforms the baselines across these variations.
 We also conducted experiments to show how perplexity changes with varying $\rho$, and the results are presented in Table 17.
@@
 The results in Table 14 show that there is no significant difference between random and structured block selection.
-5.2 Fine-tuning experiments
-Table 4: Evaluating FRUGAL for memory-efficient fine-tuning RoBERTa-Base on GLUE benchmark. Results represent the mean and standard deviation across 3 independent runs. Upper $\uparrow$ is better.
-Method Modules Rank CoLA STS-B MRPC RTE SST2 MNLI QNLI QQP Avg
-LoRA QV 8 63.8$\pm$.6 90.9$\pm$.1 89.1$\pm$.4 79.2$\pm$1.1 94.8$\pm$.2 87.6$\pm$.2 93.1$\pm$.1 90.6$\pm$.0 86.1
-GaLore All 8 60.0$\pm$.2 90.8$\pm$.1 89.0$\pm$.7 79.7$\pm$.9 94.9$\pm$.5 87.6$\pm$.1 93.3$\pm$.1 91.1$\pm$.1 85.8
-GaLore QV 8 56.1$\pm$.8 90.8$\pm$.2 88.1$\pm$.3 74.7$\pm$1.9 94.3$\pm$.1 86.6$\pm$.1 92.6$\pm$.1 89.4$\pm$.1 84.1
-FRUGAL QV 8 64.5$\pm$.7 91.1$\pm$.1 89.2$\pm$.3 82.4$\pm$.9 94.8$\pm$.2 87.4$\pm$.1 92.8$\pm$.1 91.4$\pm$.1 86.7
-FRUGAL None 0 64.8$\pm$.5 91.1$\pm$.1 89.1$\pm$.3 81.6$\pm$.6 94.9$\pm$.2 87.3$\pm$.1 92.8$\pm$.1 91.3$\pm$.1 86.6
-We evaluated the performance of our framework in the context of memory-efficient fine-tuning using the GLUE benchmark (Wang, 2018), a widely-used collection of tasks for evaluating language models.
+These experimental results validate that our framework’s superiority is resilient to hyperparameter variations.
+7 Fine-tuning experiments
+7.1 Fine-tuning RoBERTa on GLUE
+We evaluated the performance of our framework in memory-efficient fine-tuning using the GLUE benchmark (Wang, 2018), a widely-used collection of tasks for evaluating language models.
 Following the approach from Zhao et al.
@@
 We adhered to the setup described in LoRA, where low-rank updates of rank 8 were applied only to the Q and V matrices.
-For a detailed description of the experimental setup, see Section A.2.
-However, this comparison required a minor modification to FRUGAL compared to the pre-training phase.
-Instead of selecting active parameters blockwise, we opted for columnwise selection in each matrix.
-This adjustment was necessary to ensure a fair comparison within a similar memory budget, as the number of trainable parameters in LoRA with rank 8 is approximately $2.5$ times fewer than the number of trainable parameters in any RoBERTa matrix.
-This transition from blockwise to columnwise selection allowed us to maintain comparable memory usage across methods.
+See detailed description in Section A.2.
+For this experiment we opted for columnwise selection of active parameters.
+This transition from blockwise to columnwise selection was necessary to maintain comparable memory usage across methods, as the number of trainable parameters in LoRA with rank 8 is approximately $2.5$ times fewer than the number of parameters in any RoBERTa matrix.
 For the same reason, we did not include comparisons with BAdam (Luo et al., 2024) in this setup.
@@
 The results demonstrate that FRUGAL significantly outperforms GaLore and shows comparable results to LoRA.
-As in Section 5.1.1, we conducted additional experiments with FRUGAL using $\rho=0.0$.
-In this setup, only the classification head is trained using Adam, while the embedding parameters remain frozen, and the remaining parameters are trained using signSGD.
+As in Section 6.1, we conducted additional experiments with FRUGAL using $\rho=0.0$.
+In this setup, only the classification head is trained using AdamW, while the embedding parameters remain frozen, and the remaining parameters are trained using signSGD.
 The results demonstrate that this training approach barely compromises performance compared to FRUGAL with rank 8, and still outperforms GaLore.
-Similar to our findings in Section 5.1.1, we observe that the classification head parameters are particularly sensitive to the choice of optimizer, which can be seen in Table 13 where the model’s performance significantly deteriorates when using signSGD for classification head optimization.
+Similar to our findings in Section 6.1, we observe that the classification head parameters are particularly sensitive to the choice of optimizer, which can be seen in Table 19 where the model’s performance significantly deteriorates when using signSGD for classification head optimization.
+7.2 Fine-tuning LLaMA on commonsense reasoning
+In addition to our experiments with RoBERTa, we conducted experiments on LLM fine-tuning to evaluate our framework’s performance in this practically important area.
+To assess this capability, we chose LLaMA 3.1-8B (Dubey et al., 2024) and commonsense reasoning benchmark, as this domain represents a fundamental capability requiring both factual knowledge and logical inference.
+This benchmark includes 8 subtasks each containing its own training and test splits.
+Following the setup from Hu et al.
+(2023) we train the model on a single combined Commonsense170K dataset (Hu et al., 2023).
+We refer readers to the original paper for detailed descriptions of these tasks and the construction of the Commonsense170K dataset.
+Following the experimental protocol from Hu et al.
+(2023), we apply memory-efficient methods to the same parameter subsets: the Q, K, V, Up, and Down projection matrices.
+We used the same hyperparameter configuration as in the original work, except for the learning rate, which we varied across [5e-6, 1e-5, 2e-5, 5e-5, 1e-4, 2e-4] for each algorithm to ensure optimal performance.
+The results in Table 7 show that FRUGAL again slightly outperforms both LoRA and GaLore in the average accuracy across all 8 tasks.
+Remarkably, our method attains this advantage even with $\rho=0$ (effectively signSGD), requiring zero memory for optimizer state storage while delivering better accuracy.
 8 Conclusion
@@
 We prove theoretical convergence guarantees for our framework with SGDM serving as the state-full algorithm and SGD as the state-free algorithm.
-In experiments involving pre-training and fine-tuning of language models, FRUGAL outperforms other approaches while using the same or smaller memory.
-7 Acknowledgments
-The work was done in the Laboratory of Federated Learning Problems of the ISP RAS (Supported by Grant App.
-No.
-2 to Agreement No.
-075-03-2024-214).
+In experiments involving pre-training and fine-tuning of language models, FRUGAL outperforms other approaches.
+Acknowledgements
+The work on the final version was conducted at Moscow Institute of Physics and Technology and was supported by the Ministry of Economic Development of the Russian Federation (code 25-139-66879-1-0003).
 The work was partially conducted while Philip Zmushko was a visiting research assistant in Mohamed bin Zayed University of Artificial Intelligence (MBZUAI).
-Appendix A Additional experimental details
+Impact Statement
+Our work makes large-scale model training more accessible to the broader research community by reducing memory requirements.
+This reduction in computational demands not only decreases the financial cost of training but also potentially lowers the environmental impact.
+The proposed methods enable researchers with limited resources to participate in large-scale machine learning research, potentially accelerating progress in the field through wider community involvement.
+Appendix A Experimental setups
 This section describes the main setups used in the experiments and presents additional experiments.
@@
 For GaLore, this parameter is equal to $\rho=r/h$, where $r$ is the projection rank, and $h$ is the hidden size of the model.
-For RandK projection, this parameter can be expressed as $1-s$, where $s$ means sparsity.
-For BAdam and FRUGAL with the blockwise update, this parameter denotes the ratio of the number of active blocks $a_{\text{block}}$ to the total number of blocks $p$, i.e., $\rho=a_{\text{block}}/p$.
-When using FRUGAL with the column-wise update, as in Section 5.2, $\rho$ is equal to the ratio of the number of active columns $a_{\text{column}}$ to their total number $h$, i.e., $\rho=a_{\text{column}}/h$.
-A.1 Pre-training
-Setup.
-We adopt a LLaMA-based architecture with RMSNorm and SwiGLU (Wang, 2018) activations on the C4 dataset.
+For the RandK projection, this parameter can be expressed as $1-s$, where $s$ means sparsity.
+For BAdam and FRUGAL with the blockwise update, this parameter denotes the ratio of the number of active blocks $a_{\text{block}}$ to the total number of blocks $p$, that is, $\rho=a_{\text{block}}/p$.
+When using FRUGAL with the column-wise update, as in Section 7, $\rho$ is equal to the ratio of the number of active columns $a_{\text{column}}$ to their total number $h$, i.e., $\rho=a_{\text{column}}/h$.
+A.1 Pre-training setup
+We adopt a LLaMA-based architecture with RMSNorm (Zhang & Sennrich, 2019) and SwiGLU (Shazeer, 2020) activations on the C4 dataset.
 Following Zhao et al.
@@
 We used T5 tokenizer, since it also was trained on C4 with dictionary size equal to 32k.
-Additionally, the update frequency $T$ is set to 200, and the density $\rho=r/h$ is $0.25$ unless stated otherwise.
+The update frequency $T$ is set to 200.
+Since, unlike GaLore, we consider not only matrix projections, we decided to generalize the concept of rank $r$.
+Instead, we use density $\rho$, which represents the proportion of Linear layer parameters in the state-full subspace.
+Thus, for SVD-like projection as in GaLore, the density equals $\rho=r/h$, where $h$ denotes the hidden dimension of the model.
 We also should point out that similarly to Zhao et al.
-(2024a), we keep Embeddings, RMSNorms (Zhang & Sennrich, 2019), and Logits in the state-full subspace throughout the training and don’t reset the optimizer state for them.
+(2024a), we keep Embeddings, RMSNorms, and Output layer in the state-full subspace throughout the training and don’t reset the optimizer state for them.
 We used standard Adam hyperparameters: $\beta_{1}=0.9,\beta_{2}=0.999,\varepsilon=1e-8$.
-For all the methods except GaLore, we selected the learning rate equal to the optimal learning rate for Adam, which we determined through a grid search among values $[1e-4,3e-4,1e-3,3e-3]$.
+For all methods except GaLore, we selected the learning rate equal to the optimal learning rate for Adam, which we determined through a grid search among values $[1e-4,3e-4,1e-3,3e-3]$.
 FRUGAL’s learning rate for the state-free optimizer was set equal to that for the state-full optimizer for simplicity and ease of tuning.
-For a fair comparison with GaLore (Zhao et al., 2024a), we conducted experiments with two learning rate values: 1) the one specified by the authors in the original paper, and 2) the optimal learning rate for Adam, as used for other methods.
+For a fair comparison with GaLore (Zhao et al., 2024a), we conducted experiments with two learning rate values: 1) the one specified by the authors in the original paper and 2) the optimal learning rate for Adam, as used for other methods.
 We did this because the learning rate in the original paper could have been optimized for a different number of iterations.
@@
 To verify that our results are not sensitive to the choice of scheduler, we repeated the experiments for LLaMA-130M with other schedulers.
-Results for constant with warm-up and cosine (one cycle) with warm-up schedulers can be found in Tables 6 and 6.
+The results for constant with warm-up and cosine (one cycle) with warm-up schedulers can be found in Tables 16 and 16.
+For pre-training GPT-2 124M (Radford et al., 2019) we followed the setup described above except for the tokenizer.
+We utilized the GPT-2 original tokenizer, with 50257 vocabulary size.
+The results are presented in Table 12.
+Table 8: Comparison of validation perplexity and memory estimation for various optimization methods across LLaMA model scales trained on C4 with $\bm{\beta_{2}=0.95}$.
+We also indicate the additional memory overhead introduced by the optimization algorithm.
+The values are calculated assuming that each float value occupies 4 bytes (float32).
+$\rho$ denotes the proportion of the Linear layer parameters in the state-full subspace.
+Note that Embeddings, RMSNorms, and Output layer are always trained with AdamW.
+60M 130M 350M
+AdamW 23.51 (0.43G) 18.32 (1.00G) 14.57 (2.74G)
+GaLore, $\rho=0.25$ 26.66 (0.30G) 21.03 (0.54G) 16.79 (1.10G)
+BAdam, $\rho=0.25$ 25.40 (0.29G) 20.17 (0.52G) 16.54 (1.05G)
+FRUGAL, $\rho=0.25$ 24.07 (0.29G) 18.79 (0.52G) 14.96 (1.05G)
+FRUGAL, $\rho=0.0$ 24.54 (0.24G) 19.11 (0.37G) 15.20 (0.49G)
+Training tokens 20B 20B 24B
+Number of iterations 200k 200k 240k
+Table 9: Perplexity of LLaMA-130M models pre-trained on C4 using pure bfloat16 format both for model weights and optimizer statistics.
+Method 100k iterations
+Adam 21.88
+GaLore, $\rho=0.25$ 24.19
+BAdam, $\rho=0.25$ 25.03
+FRUGAL, $\rho=0.25$ 23.17
+FRUGAL, $\rho=0.0$ 22.64
+Table 10: Perplexity of LLaMA-130M models pre-trained on C4 for 200k steps with different state-free optimizers for FRUGAL.
+Method State-free Validation
+optimizer perplexity
+Adam 18.13
+FRUGAL, $\rho=0.25$ signSGD 18.60
+FRUGAL, $\rho=0.25$ SGD 19.11
+Table 11: Perplexity of LLaMA-130M models pre-trained on C4 with Lion as state-full optimizer for 200k steps.
+Method 200k
+Adam 18.13
+Lion 18.55
+GaLore (+ Lion), $\rho=0.25$ 21.65
+FRUGAL (+ Lion), $\rho=0.25$ 18.89
+Table 12: Validation perplexity of GPT-2 124M model pre-trained on C4 for 200k steps with various optimization methods.
+Method Validation perplexity
+Adam 21.94
+GaLore, $\rho=0.25$ 25.84
+BAdam, $\rho=0.25$ 25.43
+FRUGAL, $\rho=0.25$ 23.23
+FRUGAL, $\rho=0.0$ 25.04
+Table 13: Perplexity of LLaMA-130M models pre-trained on C4 for 200k iterations using FRUGAL with $\rho=1/3$ and different Block update strategy, taken from Luo et al. (2024).
+Method Perplexity
+Random 18.50
+Ascending 18.54
+Descending 18.50
+Table 14: Perplexity of LLaMA-130M models pre-trained on C4 for 200k iterations (20B tokens) using FRUGAL with $\rho=0.25$ and different update frequency $T$.
+Update frequency $T$ Perplexity
+10 18.82
+20 18.73
+50 18.69
+100 18.65
+200 18.60
+500 18.60
+1000 18.61
 Table 15: Perplexity of LLaMA-130M models pre-trained on C4 using constant scheduler with warm-up at various training iterations.
@@
 FRUGAL, $\rho=0.0$ 19.83 18.34
-Table 7: Perplexity of LLaMA-130M models pre-trained on C4 using pure bfloat16 format both for model weights and optimizer statistics.
-Method 100k iterations
-Adam 21.88
-GaLore, $\rho=0.25$ 24.19
-BAdam, $\rho=0.25$ 25.03
-FRUGAL, $\rho=0.25$ 23.17
-FRUGAL, $\rho=0.0$ 22.64
-Table 8: Perplexity of LLaMA-130M models pre-trained on C4 for 20k iterations (2.1B tokens) using SGD and signSGD with different learning rates.
-$\infty$ means that run diverged.
-LR stands for learning rate.
-SGD signSGD
-LR Perplexity LR Perplexity
-0.1 184.83 3e-4 40.22
-0.3 91.23 1e-3 41.18
-1.0 $\infty$ 3e-3 109.32
-Table 9: Perplexity of LLaMA-130M models pre-trained on C4 for 200k iterations using FRUGAL with $\rho=1/3$ and different Block update strategy, taken from Luo et al. (2024).
-Method Perplexity
-Random 18.50
-Ascending 18.54
-Descending 18.50
-Table 10: Perplexity of LLaMA-130M models pre-trained on C4 for 200k iterations (20B tokens) using FRUGAL with $\rho=0.25$ and different update frequency $T$.
-Update frequency $T$ Perplexity
-10 18.82
-20 18.73
-50 18.69
-100 18.65
-200 18.60
-500 18.60
-1000 18.61
 Table 17: Perplexity of LLaMA-130M models pre-trained on C4 for 200k iterations (20B tokens) using FRUGAL with different density $\rho$.
@@
 Perplexity 18.13 18.40 18.50 18.63 18.71 18.80 18.90 33.22
-A.2 Fine-tuning
-Setup.
+A.2 RoBERTa fine-tuning setup
 The batch size and learning rate values used for FRUGAL in the experiments from Table 6 are presented in Table 18.
@@
 None (corresponds to the fine-tuning using signSGD) 89.7 81.6 74.3
-Appendix B Memory estimation
+Appendix B Comparison with Concurrent Methods
+In this section, we present a comparison with several concurrent works — namely, Fira (Chen et al., 2024a), LDAdam (Robert et al., 2024), and Adamem (Vyas et al., 2024) — which we discovered only during the final stages of our project.
+Due to this timing, we did not have the opportunity to conduct comprehensive experimental comparisons with them as we did with other baselines in Table 2.
+B.1 Algorithmic comparison
+We begin by comparing these methods with FRUGAL from an algorithmic perspective.
+Adamem.
+Similarly to our framework, Vyas et al.
+(2024) divides the gradient into two parts: the first part is a projection of the gradient onto the top SVD subspace, while the second part is the residual outside this subspace.
+The first part is used to update momentum, which is then fed into Adafactor’s preconditioner Shazeer & Stern (2018), while the second part is fed directly into a one-sided Adafactor preconditioner Vyas et al.
+(2024).
+The outputs of these preconditioners are then used for the final parameter update.
+Although the core idea of splitting the gradient into two orthogonal subspaces is analogous to Algorithm 1, Adamem implements only one possible variant for choosing subspace and for updates of each component.
+Thus, Adamem represents a special case of FRUGAL: with SVD-based projection, Adafactor with momentum as the state-full optimizer, and one-sided Adafactor as the state-free optimizer.
+Fira.
+Chen et al.
+(2024a) also apply the idea of decomposing the full-rank gradient into a low-rank subspace gradient and a residual gradient.
+Similarly to GaLore, the low-rank part is obtained through SVD and passed through an AdamW step; however, unlike GaLore, the residual part is not discarded, but is also used in the update in an SGD-like format without using additional optimizer state.
+An important feature that Fira introduces is norm-based scaling: this technique adaptively scales the per-column learning rate for the residual part by $\frac{\|\psi_{t}(R_{t})\|}{\|R_{t}\|}$ where $\psi_{t}$ denotes the AdamW update rule.
+In addition, for enhanced training stability, Fira proposes replacing standard gradient clipping with a norm-growth limiter, which transforms abrupt gradient spikes into gradual, smooth increases.
+Thus, the key difference between Fira and FRUGAL is that the state-free update sophisticatedly utilizes information from the state-full update.
+However, Fira has several limitations.
+First, Fira relies on computing SVD for projection onto the low-rank subspace, which introduces additional memory and computational overhead, with the computational burden becoming more pronounced as model size increases (see Sections 4 and C.
+Moreover, it is not readily obvious how to propose an alternative to SVD to eliminate this overhead, because, for example, block projection from BAdam (Luo et al., 2024) cannot work with norm-based scaling (Chen et al., 2024a) by construction.
+Second, following GaLore, Fira continues using the old optimizer state after updating the projection, which should be suboptimal (see Appendix D).
+Furthermore, according to the original paper Chen et al.
+(2024a), the reason for instability in standard gradient clipping may be precisely the moments of low-rank subspace transitions.
+We also would like to note separately that neither Fira (Chen et al., 2024a) nor Adamem (Vyas et al., 2024) provide theoretical convergence proofs.
+In contrast, we provide a proof that recovers the best known convergence rate under conventional assumptions.
+LDAdam.
+Notably, unlike Fira and Adamem, Robert et al.
+(2024) identify the addition of momentum and gradient from different subspaces as mathematically incorrect (similarly to Appendix D).
+To resolve this inconsistency,  Robert et al.
+(2024), propose to reproject the previous optimizer state (similarly to Algorithm 1).
+Interestingly, LDAdam differs from FRUGAL, Adamem, and Fira in that each individual step remains low-rank, but unlike GaLore and BAdam, the remaining information is not discarded but preserved in an error feedback buffer, thus allowing LDAdam to have convergence guarantees and demonstrate high performance.
+However, LDAdam also has a drawback.
+The method requires updating the projection at every step; and although Robert et al.
+(2024) proposed replacing the expensive SVD decomposition with significantly cheaper block power iteration, the necessity of performing this operation at every step causes a slowdown of $\geq 15\%$.
+B.2 Experimental comparison
+We also present a preliminary experimental comparison of FRUGAL with Adamem (Vyas et al., 2024), Fira (Chen et al., 2024a), and LDAdam (Robert et al., 2024).
+Adamem.
+We compared FRUGAL with our reimplementation of Adamem (Vyas et al., 2024) on pre-training LLaMA model of 3 sizes — 60M, 130M, and 350M — in the same setup as in Sections 6.1 and 2.
+Unique Adamem hyperparameters were taken from Vyas et al.
+(2024).
+The results presented in Table 20 show that while Adamem performs significantly better than GaLore due to utilizing information from the gradient residual, it still falls slightly short of FRUGAL.
+However, at this point, we are not ready to claim whether this is a consequence of a less favorable choice of state-full and state-free update rules or potentially suboptimal hyperparameter settings for our experimental setup.
+Table 20: Comparison of validation perplexity for AdamW, FRUGAL and AdaMeM across LLaMA model scales trained on C4.
+60M 130M 350M
+AdamW 22.73 18.13 14.43
+AdaMeM, $\rho=0.25$ 23.81 18.99 15.10
+FRUGAL, $\rho=0.25$ 23.59 18.60 14.79
+FRUGAL, $\rho=0.0$ 24.06 18.90 15.03
+Training tokens 20B 20B 24B
+Number of iterations 200k 200k 240k
+Fira and LDAdam
+For experiments with Fira (Chen et al., 2024a) and LDAdam (Robert et al., 2024), we deviated from the setup in Appendices A and 2.
+This modification was necessary since our main setup following GaLore (Zhao et al., 2024a) does not use gradient clipping, while its counterpart—the norm-growth limiter—is critically important for Fira’s performance.
+Additionally, we observed that Fira significantly benefits from using weight decay, so for the experiments in Table 21 we enabled both gradient clipping (norm-growth limiter for Fira) and weight decay.
+The results show that all three methods achieve performance very close to AdamW.
+Surprisingly, LDAdam even outperforms the full-rank AdamW baseline for the 350M model (which, however, may be a result of insufficient hyperparameter grid search).
+Fira performs slightly better than FRUGAL.
+However, we would like to note that these improvements are valid in terms of the number of iterations, and not in terms of wall clock time, where both Fira and LDAdam introduce noticeable time overhead.
+In the same table, we report the approximate wall-clock slowdown of the methods compared to plain AdamW.
+We provide approximate values taken from the original papers because the actual slowdown can vary significantly depending on the setup (number of workers in training, GPU type, etc.).
+One can observe that the slowdown due to the need to perform additional heavyweight operations during the optimizer step can negate the advantages gained from a more carefully designed update rule.
+Thus, practitioners should choose the algorithm for their specific use case.
+Table 21: Comparison of validation perplexity for AdamW, FRUGAL  Fira and LDAdam across LLaMA model scales trained on C4.
+We also report approximate slowdown comparing to AdamW (taken from original papers).
+Approximate Slowdown 130M 350M
+AdamW - 17.52 13.81
+Fira, $\rho=0.25$ $10\%$ 17.58 13.74
+LDAdam, $\rho=0.25$ $15\%$ 17.54 13.54
+FRUGAL, $\rho=0.25$ 0% 17.58 13.99
+Appendix C Memory estimation
 In this section, we will examine memory requirements for different projection types using the LLaMA-like architecture as an example and show that RandK, column-wise, and blockwise projections result in approximately the same amount of additional memory for a given density value $\rho$ Appendix A.
@@
 Recall that we follow the setup from Zhao et al.
-(2024a), where Embeddings, RMSNorms, and Logits remain in the state-full subspace throughout the training, so the projection does not interact with them, and they give the same memory overhead for all projection methods.
+(2024a), where Embeddings, RMSNorms, and Output layer remain in the state-full subspace throughout the training, so the projection does not interact with them, and they give the same memory overhead for all projection methods.
 Let the number of parameters in the remaining projectable parameters be $P$.
@@
 Therefore, for our method, we settled on blockwise projection.
-Appendix C Optimizer state management
+Appendix D Optimizer state management
 In this section, we would like to propose some modifications to the GaLore algorithm. These modifications are also used in our framework as SVD projection.
@@
 As can be seen, the variant with state projection converges much faster.
-Appendix D LLaMA 7B pre-training results.
-In this section, we present the results of pre-training a LLaMA 7b model on the C4 dataset for 120k iterations on 12B tokens.
-See results in Table 14.
-We conducted the training in pure bfloat16 with the density $\rho=0.0$.
-We used learning rate $0.0005$ for state-full optimizer and $0.00015$ for state-free optimizer.
-However, unlike Zhao et al.
-(2024a), we didn’t use Adam8bit for state-full parameters but rather Adam, so it may not be an entirely fair comparison.
-Nevertheless, the results show that FRUGAL has the potential for scaling up to 7B parameter models.
-Table 14: Pre-training LLaMA 7B on C4 dataset for 120K steps.
-Validation perplexity and memory estimate are reported.
-We also indicate the additional memory overhead introduced by the optimization algorithm.
-40K 80K 120K
-8-bit Adam 18.09 15.47 14.83
-8-bit GaLore 17.94 15.39 14.95
-FRUGAL, $\rho=0.0$ 17.56 14.50 13.49
-Tokens (B) 5.2 10.5 15.7
 Appendix E Convergence Theory
 Firstly, we provide ommited definition of $L$-smooth function.
-Definition 1.
+Definition E.1.
 We say that $f:\operatorname*{\mathbb{R}^{d}}\rightarrow\mathbb{R}$ is $L-$smooth with $L\geq 0$, if it is differentiable and satisfies
 $f(y)\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{L}{2}\|y-x\|^{2},\forall x,y\in\operatorname*{\mathbb{R}^{d}}.$
-Below, we provide an equivalent formulation of Algorithm 2 that enables us to use the proof of the similar structure to SGDM momentum analyis of Liu et al. (2020).
-Algorithm 3 FRUGAL(SGDM, SGD): Equivalent to Algorithm 2 for constant step isze
-Input: momentum weight $\beta\in[0,1)$, initialization $x^{1}\in\operatorname*{\mathbb{R}^{d}}$ and $m^{0}=0$, step sizes $\{\alpha_{k}:=\alpha>0\}_{k=1}^{K}$, momentum set $J_{k}\subset[d]$ for $k=1,2\ldots$.
+Below, we provide an equivalent formulation of Algorithm 2 that enables us to use the proof of the similar structure to SGDM momentum analyis of (Liu et al., 2020).
+Algorithm 3 FRUGAL(SGDM, SGD): Equivalent to Algorithm 2 for constant step size
+Input: momentum weight $\beta\in[0,1)$, initialization $x^{1}\in\operatorname*{\mathbb{R}^{d}}$ and $m^{0}=0$, step sizes $\{\alpha_{k}:=\alpha>0\}_{k=1}^{K}$, momentum set $J_{k}\subset[d]$ for $k=1,2\ldots$
 1:  for $k=1,2,\ldots$ do
-2:     Compute stochastic gradient $\tilde{g}^{k}\leftarrow\nabla f_{\zeta^{k}}(x^{k})$;
-3:     Update momentum vector $\tilde{m}_{j}^{k}\leftarrow(1-\beta)\tilde{g}_{j}^{k}+\beta\begin{cases}\tilde{m}_{j}^{k-1}&\text{if }j\in J_{k},\\ 0&\text{otherwise;}\end{cases}$
-4:     Update iterate $x^{k+\nicefrac{{1}}{{2}}}\leftarrow x^{k}-\alpha\tilde{m}^{k}$;
-5:     $x_{j}^{k+1}\leftarrow\begin{cases}\frac{x_{j}^{k+\nicefrac{{1}}{{2}}}}{1-\beta}-\frac{\beta x^{k}_{j}}{1-\beta}&\text{if }j\notin J_{k+1},\\ x_{j}^{k+\nicefrac{{1}}{{2}}}&\text{otherwise;}\end{cases}$
+2:     Compute stochastic gradient $\tilde{g}^{k}\leftarrow\nabla f_{\zeta^{k}}(x^{k})$
+3:     Update momentum vector $\tilde{m}_{j}^{k}\leftarrow(1-\beta)\tilde{g}_{j}^{k}+\beta\begin{cases}\tilde{m}_{j}^{k-1}&\text{if }j\in J_{k},\\ 0&\text{otherwise}\end{cases}$
+4:     Update iterate $x^{k+\nicefrac{{1}}{{2}}}\leftarrow x^{k}-\alpha\tilde{m}^{k}$
+5:     $x_{j}^{k+1}\leftarrow\begin{cases}\frac{x_{j}^{k+\nicefrac{{1}}{{2}}}}{1-\beta}-\frac{\beta x^{k}_{j}}{1-\beta}&\text{if }j\notin J_{k+1},\\ x_{j}^{k+\nicefrac{{1}}{{2}}}&\text{otherwise}\end{cases}$
 6:  end for
@@
 Using this notation, we proceed with two lemmas, one showing variance reduction effect of momentum, the other boundess of momentum bias.
-Lemma 1.
-Under Assumption 1, the update vector $\tilde{m}^{k}$ in Algorithm 3 satisfies
+Lemma E.2.
+Under Assumption 5.1, the update vector $\tilde{m}^{k}$ in Algorithm 3 satisfies
 $\mathbb{E}\left[\left\|\tilde{m}^{k}-m^{k}\right\|^{2}\right]$ $\leq\frac{1-\beta}{1+\beta}\sigma^{2}.$
@@
 $\leq(1-\beta)^{2}\sum_{j\in[d]}\mathbb{E}\left[\left\|\sum_{i=t^{k}_{j}}^{k}\beta^{k-i}(\tilde{g}_{j}^{i}-g_{j}^{i})\right\|^{2}\right].$
-Moreover, since $\zeta^{1},\zeta^{2},...,\zeta^{k}$ are independent random variables (item 3 of Assumption 1), we can use conditional expectation to show that $\mathbb{E}\left[(\tilde{g}_{j}^{i_{1}}-g_{j}^{i_{1}})(\tilde{g}_{j}^{i_{2}}-g_{j}^{i_{2}})\right]=0$ for $i_{1}\neq i_{2}$.
+Moreover, since $\zeta^{1},\zeta^{2},...,\zeta^{k}$ are independent random variables (item 3 of Assumption 5.1), we can use conditional expectation to show that $\mathbb{E}\left[(\tilde{g}_{j}^{i_{1}}-g_{j}^{i_{1}})(\tilde{g}_{j}^{i_{2}}-g_{j}^{i_{2}})\right]=0$ for $i_{1}\neq i_{2}$.
 Therefore,
@@
 $\leq\frac{1-\beta}{1+\beta}\sum_{j\in[d]}\sigma_{j}^{2}=\frac{1-\beta}{1+\beta}\sigma^{2}.$
-Lemma 2.
-Under Assumption 1, the update vector $\tilde{m}^{k}$ in Algorithm 3 further satisfies
+Lemma E.3.
+Under Assumption 5.1, the update vector $\tilde{m}^{k}$ in Algorithm 3 further satisfies
 $\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]\leq p^{k}_{\max}\mathbb{E}\left[\sum_{i=1}^{k-1}a_{k,i}\|x^{i+1}-x^{i}\|^{2}\right],$
@@
 where we applied Cauchy-Schwarz to the first inequality.
-By applying triangle inequality and the smoothness of $f$ (item 1 in Assumption 1), we further have
+By applying triangle inequality and the smoothness of $f$ (item 1 in Assumption 5.1), we further have
 $\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]$ $\leq(1-\beta)p^{k}_{\max}\mathbb{E}\left[\sum_{i=1}^{k}\beta^{k-i}(k-i)\sum_{l=i}^{k-1}\|g^{l+1}-g^{l}\|^{2}\right]$
@@
 $a_{k,i}=L^{2}\beta^{k-i}\left(k-i+\frac{\beta}{1-\beta}\right).$
-From Lemma 2, we know that the distance of the non-stochastic momentum from $g^{k}$ is bounded by the weighted sum of past successive iterate differences.
+From Lemma E.3, we know that the distance of the non-stochastic momentum from $g^{k}$ is bounded by the weighted sum of past successive iterate differences.
 Furthermore, the coefficients $a_{k,i}$ decays exponentially in $\beta$.
@@
 which behaves more like an SGD iterate, although the stochastic gradient $\tilde{g}^{k}$ is not taken at $z^{k}$.
-Lemma 3.
+Lemma E.4.
 Let $x^{k}$’s be iterates of Algorithm 3, then $z^{k}$ defined in equation 10 satisfies
@@
 Before procceding with the main convergence theory, we require one more proposition that shows descent in objective value.
-Proposition 1.
-Take Assumption 1. Then, for $z^{k}$ defined in equation 10, we have
+Proposition E.5.
+Take Assumption 5.1. Then, for $z^{k}$ defined in equation 10, we have
 $\begin{split}\mathbb{E}[f(z^{k+1})]&\leq\mathbb{E}[f(z^{k})]+\left(-\alpha+\frac{1+\beta^{2}}{1-\beta}L\alpha^{2}+\frac{1}{2}L\alpha^{2}\right)\mathbb{E}[\|g^{k}\|^{2}]\\ &\quad+\left(\frac{\beta^{2}}{2(1+\beta)}+\frac{1}{2}\right)L\alpha^{2}\sigma^{2}+\frac{L\alpha^{2}}{1-\beta}\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right].\end{split}$ (11)
@@
 $\begin{split}\mathbb{E}_{\zeta^{k}}[f(z^{k+1})]&\leq f(z^{k})+\mathbb{E}_{\zeta^{k}}[\langle\nabla f(z^{k}),z^{k+1}-z^{k}\rangle]+\frac{L}{2}\mathbb{E}_{\zeta^{k}}[\|z^{k+1}-z^{k}\|^{2}]\\ &=f(z^{k})+\mathbb{E}_{\zeta^{k}}[\langle\nabla f(z^{k}),-\alpha\tilde{g}^{k}\rangle]+\frac{L\alpha^{2}}{2}\mathbb{E}_{\zeta^{k}}[\|\tilde{g}^{k}\|^{2}],\end{split}$ (12)
-where we have applied Lemma 3 in the second step.
+where we have applied Lemma E.4 in the second step.
 For the inner product term, we can take full expectation $\mathbb{E}=\mathbb{E}_{\zeta^{1}}...\mathbb{E}_{\zeta^{k}}$ to get
@@
 $\begin{split}\mathbb{E}[f(z^{k+1})]&\leq\mathbb{E}[f(z^{k})]+\alpha^{3}\frac{\rho_{0}}{2}L^{2}(\frac{\beta}{1-\beta})^{2}\mathbb{E}\left[\sum_{j\in J_{k}}\|\tilde{m}_{j}^{k-1}\|^{2}\right]\\ &\quad+(\alpha\frac{1}{2\rho_{0}}-\alpha)\mathbb{E}[\|g^{k}\|^{2}]+\frac{L\alpha^{2}}{2}\mathbb{E}[\|\tilde{g}^{k}\|^{2}].\end{split}$ (13)
-Let $k_{j}=k-t^{k-1}_{j}+1$. Then, from Lemma 1 we know that
+Let $k_{j}=k-t^{k-1}_{j}+1$. Then, from Lemma E.2 we know that
 $\begin{split}\mathbb{E}\left[\sum_{j\in J_{k}}\|\tilde{m}_{j}^{k-1}\|^{2}\right]&\leq 2\mathbb{E}\left[\sum_{j\in J_{k}}\|\tilde{m}_{j}^{k-1}-m_{j}^{k-1}\|^{2}\right]+2\mathbb{E}\left[\sum_{j\in J_{k}}\|m_{j}^{k-1}\|^{2}\right]\\ &\leq 2\frac{1-\beta}{1+\beta}\mathbb{E}\left[\sum_{j\in J_{k}}\sigma_{j}^{2}+2\sum_{j\in J_{k}}\|m_{j}^{k-1}\|^{2}\right]\\ \mathbb{E}\left[\sum_{j\in J_{k}}\|m_{j}^{k-1}\|^{2}\right]&=\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{(k-1)_{j}})^{2}\left\|\frac{m_{j}^{k-1}}{(1-\beta^{(k-1)_{j}})}\right\|^{2}\right]\\ &\leq 2\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{(k-1)_{j}})^{2}\left\|\frac{m_{j}^{k-1}}{(1-\beta^{(k-1)_{j}})}-g_{j}^{k}\right\|^{2}\right]+2\mathbb{E}\left[\sum_{j\in J_{k}}\left\|g_{j}^{k}\right\|^{2}\right]\\ \mathbb{E}\left[\|\tilde{g}^{k}\|^{2}\right]&\leq\sigma^{2}+\mathbb{E}[\|g^{k}\|^{2}].\end{split}$ (14)
@@
 Firstly, by combining results from prior section, we can bound our Lyapunov function $L^{k}$ defined in equation 9.
-Proposition 2.
-Let Assumption 1 hold and $\alpha\leq\frac{1-\beta}{2\sqrt{2}L\sqrt{p^{k}_{\max}}\sqrt{\beta+\beta^{2}}}$ in Algorithm 3. Let $\{c_{i}\}_{i=1}^{\infty}$ in equation 9 be defined by
+Proposition E.6.
+Let Assumption 5.1 hold and $\alpha\leq\frac{1-\beta}{2\sqrt{2}L\sqrt{p^{k}_{\max}}\sqrt{\beta+\beta^{2}}}$ in Algorithm 3. Let $\{c_{i}\}_{i=1}^{\infty}$ in equation 9 be defined by
 $c_{1}$ $=\frac{\frac{\beta+\beta^{2}}{(1-\beta)^{3}}L^{3}\alpha^{2}}{1-4\alpha^{2}\frac{\beta+\beta^{2}}{(1-\beta)^{2}}L^{2}},\quad\quad c_{i+1}=c_{i}-\left(4c_{1}\alpha^{2}+\frac{L\alpha^{2}}{1-\beta}\right)\beta^{i}(i+\frac{\beta}{1-\beta})L^{2}\quad\text{for all $i\geq 1$.}$
@@
 In the rest of the proof, let us show that the sum of the last two terms in equation 19 is non-positive.
-First of all, by Lemma 2 we know that
+First of all, by Lemma E.3 we know that
 $\mathbb{E}\left[\sum_{j\in J_{k}}(1-\beta^{k_{j}})^{2}\left\|\frac{m_{j}^{k}}{(1-\beta^{k_{j}})}-g_{j}^{k}\right\|^{2}\right]\leq\mathbb{E}\left[p^{k}_{\max}\sum_{i=1}^{k-1}a_{k,i}\|x^{i+1}-x^{i}\|^{2}\right],$
@@
 By telescoping equation 16, we obtain the convergence bound of our proposed algorithm under nonconvex settings.
-Theorem 2.
-Let Assumption 1 hold and $\alpha^{k}=\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$. Then, the iterates of Algorithm 3 satisfy
+Theorem E.7.
+Let Assumption 5.1 hold and $\alpha^{k}=\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$. Then, the iterates of Algorithm 3 satisfy
 $\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}[\|g^{i}\|^{2}]$ $\leq\mathcal{O}\left(\frac{f(x^{1})-f^{*}}{k\alpha}+L\alpha\sigma^{2}\left(1+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{(1-\beta)}\right)\right),$ (21)
@@
 $\begin{split}\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}[\|g^{i}\|^{2}]&\leq\frac{2\left(f(x^{1})-f^{*}\right)}{k\alpha}+\left(\frac{2\beta^{2}+8\hat{p}^{k}_{\max}}{2(1+\beta)}+\frac{1}{2}+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{8(1-\beta)}\right)L\alpha\sigma^{2}\\ &=\mathcal{O}\left(\frac{f(x^{1})-f^{*}}{k\alpha}+L\alpha\sigma^{2}\left(1+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{(1-\beta)}\right)\right).\end{split}$
+Appendix F Limitations
+We would also like to acknowledge the limitations of this work.
+Due to computational constraints, we were unable to conduct experiments on pre-training 7B+ LLMs, which is crucial for understanding the potential of our approach when scaling.
+Furthermore, our experiments are limited to training language models, although memory-efficient optimization could also be beneficial for training diffusion models.
+Finally, there may be a better method for selecting the next state-full subspace during the training.
+We leave the exploration of more sophisticated selection strategies for future work.
+Appendix G Simplified algorithms pseudocode
+In this section we present the simplified pseudocode of FRUGAL.
+In Algorithm 4 one can find optimizer steps both for FRUGAL with SVD projection (GaLore-like (Zhao et al., 2024a)) and Block projection (BAdam-like (Luo et al., 2024)).
+1:  def svd_or_randk_step(self):
+2:    for param in self.params:
+3:      grad = param.grad
+4:      param_state = self.state[param]
+5:      # update projector if necessary
+6:      if self.step % self.update_gap == 0:
+7:        param_state["projector"] = self.update_proj(grad)
+8:      projector = param_state["projector"]
+9:      # obtain state-full grad and state-free grad
+10:      grad_full = projector.proj_down(grad)
+11:      grad_free = grad_full - projector.proj_up(grad_full)
+12:      # reset state-full optimizer state if necessary
+13:      if self.step % self.update_gap == 0:
+14:        param_state["exp_avg"] = torch.zeros_like(grad_full)
+15:        param_state["exp_avg_sq"] = torch.zeros_like(grad_full)
+16:      # state-full subspace update
+17:      self.step += 1
+18:      update_full = self.state_full_step(grad_full, param_state)
+19:      update_full = projector.proj_up(update_full)
+20:      # state-free subspace update
+21:      update_free = self.state_free_step(grad_free)
+22:      # perform resulting update
+23:      update = update_full + update_free
+24:      param.add_(update)
+25:
+26:  def block_step(self):
+27:    # change state-full and state-free blocks if necessary
+28:    if self.step % self.update_gap == 0:
+29:      indices_full = self.update_indices(indices_full)
+30:      for idx, param in enumerate(self.params):
+31:        grad = param.grad
+32:        param_state = self.state[param]
+33:        if idx in indices_full:
+34:          # reset state-full optimizer state
+35:          param_state["exp_avg"] = torch.zeros_like(grad)
+36:          param_state["exp_avg_sq"] = torch.zeros_like(grad)
+37:          param_state["full_subspace"] = True
+38:        else:
+39:          # free state-full optimizer state to save memory
+40:          param_state.clear()
+41:          param_state["full_subspace"] = False
+42:    # perform updates
+43:    for param in self.params:
+44:      grad = param.grad
+45:      param_state = self.state[param]
+46:      # choose the optimizer depending on the block type
+47:      if param_state["full_subspace"]:
+48:        update = self.state_full_step(grad, param_state)
+49:      else:
+50:        update = self.state_free_step(grad)
+51:      # perform resulting update
+52:      param.add_(update)
+Algorithm 4 FRUGAL step pseudocode, PyTorch-like
+1:  def state_full_adam_step(self, grad, param_state):
+2:    exp_avg = param_state["exp_avg"]
+3:    exp_avg_sq = param_state["exp_avg_sq"]
+4:    step = self.step
+5:    beta1, beta2 = self.betas
+6:    exp_avg.mul_(beta1).add_(grad, alpha=1.0-beta1)
+7:    exp_avg_sq.mul_(beta2).addcmul_(grad, grad, value=1.0-beta2)
+8:    denom = exp_avg_sq.sqrt()
+9:    step_size = self.lr_full
+10:    if self.correct_bias:
+11:      bias_correction1 = 1.0 - beta1 ** step
+12:      bias_correction2 = 1.0 - beta2 ** step
+13:      step_size = self.lr_full / bias_correction1
+14:      bias_correction2_sqrt = math.sqrt(bias_correction2)
+15:      denom.div_(bias_correction2_sqrt)
+16:    denom.add_(self.eps)
+17:    update_full = exp_avg / denom * (-step_size)
+18:    return update_full
+19:
+20:  def state_free_signsgd_step(self, grad):
+21:    update_free = -self.lr_free * grad.sign()
+22:    return update_free
+Algorithm 5 Examples of state-full and state-free steps for Algorithm 4